### v1

# Making Every Frame Matter: Continuous Video Understanding for Large Models via Adaptive State Modeling

Abstract.
Video understanding has become increasingly important with the rise of multi-modality applications. Understanding continuous video poses considerable challenges due to the fast expansion of streaming video, which contains multi-scale and untrimmed events. We introduce a novel system, C-VUE, to overcome these issues through adaptive state modeling. C-VUE has three key designs. The first is a long-range history modeling technique that uses a video-aware approach to retain historical video information. The second is a spatial redundancy reduction technique, which enhances the efficiency of history modeling based on temporal relations. The third is a parallel training structure that incorporates the frame-weighted loss to understand multi-scale events in long videos. Our C-VUE offers high accuracy and efficiency. It runs at speeds ¿30 FPS on typical edge devices and outperforms all baselines in accuracy. Moreover, applying C-VUE to a video foundation model as a video encoder in our case study resulted in a 0.46-point enhancement (on a 5-point scale) on the in-distribution dataset, and an improvement ranging from 1.19% to 4% on zero-shot datasets.
1. Introduction
Video understanding (madan2024foundation, ; chen2024video, ; huang2018makes, ; zolfaghari2018eco, ; wu2021towards, ; tang2023video, ) is an important and long-standing application for edge and mobile computing. It requires capturing spatio-temporal features from an input video to conduct high-level tasks, such as action localization, activity recognition, video question answering, and video captioning. These tasks are widely used in edge scenarios, including robotics, augmented reality, embedded devices (e.g., Kinect and Wii remote), and IoT devices. Due to user data privacy, responsiveness requirements, and the high cost of video transmission to the cloud, real-world video understanding is typically processed on edge devices.
Challenges. Compared to video processing tasks only relying on frame spatial features (wu2020emo, ; tong2021large, ; jiang2021flexible, ; voigtlaender2019mots, ; yi2020eagleeye, ; wu2021pecam, ), e.g., object detection, segmentation, and tracking, video understanding on edge is much challenging due to the following reasons.
(i) Video understanding needs to capture the spatio-temporal features of an event on different temporal scales (zhao2018trn, ; shao2020finegym, ). Figure 2 shows an indoor activity clipping from Charades dataset (sigurdsson2016hollywood, ). The temporal durations of activities vary greatly from around 155 frames (putting a bag) to 553 frames (making a sandwich). The challenge lies in modeling this variability effectively: shorter events demand quick, precise attention, while longer activities require retaining relevant context over extended periods, complicating feature extraction across different timescales.
(ii) Video understanding normally needs to capture meaningful features from the raw streaming videos, which are untrimmed and dominated by noise and redundancy in both spatial and temporal dimensions. As seen in Figure 2, the pixels that are related to an activity are very sparsely distributed in the video. The understanding may be distracted by the dominant unrelated pixels.
(iii) Video understanding on edge must guarantee a timely understanding of the scene while managing the resource constraints of typical edge devices. For instance, robots or embodied AI (duan2022survey, ) systems must continuously perceive and comprehend dynamic scenarios through video input to carry out reasoning (strafforello2023current, ) and task planning (gadre2022continuous, ).
Figure 1. C-VUE: continuous video understanding on streaming video. As each frame is received, C-VUE updates its hidden state and outputs the events that have occurred so far.
Figure 2. We utilize the video stream, along with annotations from the Charades dataset (sigurdsson2016hollywood, ), to illustrate our motivation. C-VUE aims to understand ”What has happened so far?” at any moment in a streaming video.
Figure 3. Typical designs to process frames of the video stream for continuous video understanding. $\texttt{E}_{i}$ signifies an event while $\texttt{f}_{i}$ depicts the frame at time $t_{i}$. The “All Frames at Once” design processes all video frames at every moment. The “Sliding Window” design incorporates only the latest $m$ frames. The “State Assistance” relies on a single frame at a time and archives previous data through a hidden state. Our adaptive state modeling is an improved state assistance design, excelling in long-range history modeling via a video-aware technique.
Current status of technology. Past works explore different DNN methods for video understanding, including 3D CNN (tran2015learning, ), RNN (e.g., LSTM) (hochreiter1997long, ), Transformer (vaswani2017attention, ), and the recent Mamba (gu2023mamba, ). However, none of the past works can mitigate all the challenges at the same time.
To capture the features of various temporal scales, a straightforward method is to input all the past frames into the model at each time step, as illustrated in Figure 3(a). However, it is certainly too costly for model inference and training. Take a Transformer model as an example, with the number of streaming frames $n$ grow, the computation cost of video understanding is $O(n^{3})$. Even with the many linear alternatives of Transformers, such as Linformer (wang2020linformer, ), RetNet (sun2023retentive, ), Mamba (gu2023mamba, ), and RWKV (peng2023rwkv, ), the cost is still $O(n^{2})$.
The most commonly used technique to control the cost is to fix the input size of the model even with the growing number of frames, such as sliding window or sampling, shown in Figure 3(b). For example, TSN (wang2016temporal, ) first divides the past video into segments and selects a random snippet from each segment for activity recognition. VideoLlama2 (damonlpsg2024videollama2, ) only samples 8 or 16 frames as input for any video length. This method has to ignore many detailed features.
There are also works (shi2015convolutional, ; huang2020efficient, ; zhou2018temporal, ), shown in Figure 3(c), that only input one current frame into a model (e.g., CNN) first to extract spatial features, and then conduct feature aggregation (e.g., through pooling (kar2017adascan, ) or LSTM (hochreiter1997long, )) among frames over time to capture temporal dependencies by hidden states. This is the most efficient method for inference and training, and all frames can be processed with no one lost.
This method, however, struggles to capture the temporal relations in long range, as the fundamental limitation discussed in (shao2020finegym, ; wang2016temporal, ). Current works’ limitations highlight the difficulty in creating a continuous video understanding method that meets the above three challenges. An ideal strategy should keep a fair balance between the accuracy of the Transformer and the efficiency of the LSTM, interpreting multi-scale events in rapidly growing videos. 1) It needs to correctly keep track of past state when forming decisions. 2) It must operate in real time on typical edge devices, importantly, without reducing processing speed as video length expands.
Key techniques. To address the challenges, we propose and implement C-VUE, a novel video encoder and the according training and inference system. C-VUE can perform real-time and prompt video understanding on streaming video in a continuous manner. That is to say, as each frame is received, the system updates its hidden state and outputs the events that have occurred so far, shown in Figure 1. Our C-VUE includes three key techniques. First, a long-range history modeling, designed to maintain the video historical context via the hidden state through a video-aware idea. Second, the temporal-relation guided spatial redundancy reduction, crafted to build a more efficient video modeling at the feature level. Third, a parallel training structure and a frame-weighted loss for the understanding of multi-scale events in the streaming video.
Evaluation results. We evaluate C-VUE across various downstream tasks, comparing it with widely-used models for video understanding, including Transformer (vaswani2017attention, ), RNN (LSTM) (hochreiter1997long, ), CNN (kar2017adascan, ), as well as the SOTA video understanding system for edge scenarios, i.e., MoviNetA0 (kondratyuk2021movinets, ) and Mvitv2 (li2022mvitv2, ). The evaluation is conducted to demonstrate five capabilities: (i) the understanding of events with various temporal lengths; (ii) prompt recognition of an event as more frames are received; (iii) generality to long-tail events; (iv) the localization and understanding of continuous input events; (v) long-range memorization.
The real-world video datasets used are FineGym (shao2020finegym, ) (sports gymnasium dataset), and Charades (sigurdsson2016hollywood, ) (daily indoor activities). We also compose more challenging datasets, e.g., concatenating all videos from the FineGym dataset into one to generate an hour-long video stream. Results show that with significantly fewer trainable parameters (ranging from 8.5k to 34.1M less than the baselines), C-VUE achieves accuracy improvements of 1.2% to 4.7% over Transformer, 2.6% to 5.9% over ConvPooling, 2% to 3% over LSTM, 14.5% to 67.4% over MoviNetA0, and 13.1% to 79.7% over MViTv2.
Regarding latency, we evaluate C-VUE on typical edge devices (NVIDIA Jetson Orin, Apple MacBook, and an Intel PC with NVIDIA 4090 GPU). It achieves ¿30FPS on all these devices, up to 20$\times$ faster than the available inference systems.
For the case study, we apply C-VUE video encoder to a video foundation model (Section 6). Specifically, we replace the visual encoder of the SOTA visual language model VideoLlama2 (damonlpsg2024videollama2, ) by C-VUE. The results demonstrate a 0.46-point improvement (on a 5-point scale) on the in-distribution dataset, and an improvement ranging from 1.19% to 4% on zero-shot datasets.
Contributions. We summarize our contribution as follows:
  • We propose a continuous video understanding system, C-VUE, based on adaptive state modeling. It promptly understands incoming video frames by maintaining historical context using a hidden state.
  • C-VUE employs long-range history modeling to utilize the fixed-size hidden state for memorizing extensive historical data, reduces spatial redundancy based on temporal relations, and adopts a parallel training structure along with a frame-weighted loss function.
  • Experiments show that C-VUE surpasses all baselines in continuous video understanding, using fewer parameters and achieving over 30 FPS on typical edge devices. A case study demonstrates that C-VUE can enhance the video foundation models on video QA tasks by serving as the video encoder.
2. Background
Continuous video understanding holds significant implications for complex human-computer interaction, embodied AI, and robots (duan2022survey, ; strafforello2023current, ; gadre2022continuous, ). While various approaches strive to achieve continuous understanding, none have solved this challenging task.
All Frames at Once. (Figure 3(a)) This method performs the video understanding on all video frames, theoretically yielding the highest task accuracy (slowfast2019fei, ; kondratyuk2021movinets, ). However, its efficiency is a concern. A basic transformer model has complexity $\mathcal{O}(n^{2})$, making the whole process have $\mathcal{O}(n^{3})$ complexity. Applying optimized architecture can decrease it to $\mathcal{O}(n^{2})$. Our experiment (Figure 4) shows that as frames increase, FPS drops to 0.05 times, a 95% decrease.
Fixed Window Size. (Figure 3(b)) This approach uses a fixed window size, like sliding window or sampling, to lower computational burden (Lin_2019_ICCV, ; Jiang_2019_ICCV, ; Li_2020_CVPR, ; li2022mvitv2, ). More efficient than the first method, its complexity ranges from $\mathcal{O}(mn)$ to $\mathcal{O}(m^{2}n)$, where $m$ equals the window size. The main challenge here is choosing an appropriate window size, as real-world events vary in duration. Small windows might miss longer events, and large ones could hinder understanding of shorter events.
State Assistance. (Figure 3(c)) By maintaining a group of hidden states, this method significantly enhances efficiency by reducing complexity to approximately $\mathcal{O}(n)$ (shi2015convolutional, ; huang2020efficient, ). However, as the input size grows, the fixed-size state struggles to remember the increasing history of inputs. Existing designs handle states in a video-unaware manner, preventing full utilization of the hidden states and the inability to model long-range history.
Figure 4. Processing speed of understanding a 100k-frame video on RTX4090. Baselines 1(A) and 2(A) represent Transformer and Convpooling-based solutions using the “All Frames at Once” design. Baselines 1(F) and 2(F) employ the “Fixed window size” design (the window size is 100). Baseline 3 denotes an LSTM-based solution leveraging “State Assistance” design.
3. C-VUE Design
3.1. Design Overview
Design Requirements. To continuously process the video stream, an ideal design should integrate the advantages of the aforementioned approaches. The first advantage is efficiency: input only the current frame instead of repeating or tracing back through previous frames (similar to the ”State Assistance” design). The second advantage is accuracy: have a comprehensive memory of historical information at all times (approximating the ”All Frames at Once” design).
Design Challenges. To achieve the ideal design, the key lies in managing the growing volume of video, including: First, maximizing the memorization of the historical video data in a bounded subspace. Second, enabling processing at the edge in real-time without any increase in processing time as the video length extends.
Design Preliminary. We base our method on a state-assistance design. The typical state-assistance design involves recurrent updates over a sequence through hidden states (sieber2024understanding, ). We take the RNN as an example to introduce the standard state-assistance design. It computes $\mathbf{y}$ through a dynamic recurrence of input signals $\mathbf{x}$ at each time step $t$ through hidden states $\mathbf{s}$. Formally,
(1) $s_{t}=\phi_{h}(\mathbf{A}s_{t-1}+\mathbf{B}x_{t}+\mathbf{b}_{h}),$
(2) $y_{t}=\phi_{o}(\mathbf{C}s_{t}+\mathbf{b}_{o}),$
where $\mathbf{A},\mathbf{B}$, and $\mathbf{C}$ denote the state matrix, input matrix, and output matrix, respectively. The $\mathbf{b}_{h}$ and $\mathbf{b}_{o}$ are the bias parameters, and the functions $\phi_{h}$ and $\phi_{o}$ are activation functions.
Subsequently, the gating mechanism is introduced to enhance the memory capability of state-assistance designs, e.g., LSTM (sherstinsky2020fundamentals, ). More specifically, gates denoted as $g^{s}_{t}$, $g^{i}_{t}$, and $g^{o}_{t}$ are introduced at each state, input, and output, respectively. At time $t$, these gates dynamically adjust the state update process and the output based on the given input. Formally,
(3) $s_{t}=\phi_{h}(g^{s}_{t}\circ\mathbf{A}s_{t-1}+g^{i}_{t}\circ\mathbf{B}x_{t}+\mathbf{b}_{h}),$
(4) $y_{t}=\phi_{o}(g^{o}_{t}\circ\mathbf{C}s_{t}+\mathbf{b}_{o}).$
This gating mechanism resembles a weighting scheme, represented as a vector with values ranging from 0 to 1, which is computed based on the input and the last output.
Limitation of RNN and Its Variants. The RNN and its gating-based variants are widely considered difficult to train and challenging to handle long-range historical data by many research works (greaves2019statistical, ; zhao2020rnn, ; miller2018stable, ). Our experimental results also indicate that the LSTM-based baseline performs poorly on hours-long videos (Table 2).
We argue that the underperformance of such methods is due to their inefficient utilization of the hidden state. Specifically, standard RNN processes input at different moments identically, i.e., fixed $A$, $B$, and $C$ in Equations 1 and 2, resulting in a lack of adaptation towards these inputs. When redundancies exist within the input data, this leads to an irrational use of the hidden state. Despite the subsequent introduction of gating mechanisms to enhance input adaptation, these designs merely adopt a weighted approach to decide how the new inputs should merge with the existing states. Such a design idea lacks flexibility and fails to fully tap into the memorization capability of the hidden state.
Our Key Designs. We introduce our C-VUE with the following three parts. First, long-range history modeling (Section 3.2). This is our core design, aiming to maximize the memorization capability of multi-scale events over long sequences. This design adopts two techniques: 1) a fully video-aware mechanism to improve the hidden state’s utilization; and 2) a scale invariance mechanism for multi-scale events understanding.
Second, temporal relation-guided spatial redundancy reduction (Section 3.3). Performing history modeling in pixel space can result in excessive system overhead and increased learning difficulty. This design enables the efficient modeling by eliminating the redundant information at the feature level with the guidance of the temporal relation.
Third, parallal training with frame-weighted loss (Section 3.4). It introduces how our C-VUE builds the training structure for continuous long video understanding. We detail the efficient parallel training for long video samples and the weighted loss design for the understanding of every frame.
3.2. Long-range History Modeling
The challenge of understanding long-range history primarily stems from two aspects, namely inter-event noise and intra-event sparsity. The inter-event noise implies that there may be numerous noise frames between two events. In the context of continuous understanding, the longer the video, the more pronounced this issue becomes. The intra-event sparsity represents the inherent feature sparsity of event information. Some events carry limited information despite their extended durations, e.g., sleeping. However, the multi-scale characteristics of events make it difficult to handle this sparsity.
To handle these issues, we propose two techniques. The first is a fully video-aware history modeling. Here, each fixed parameter is improved to a dynamically-generated parameter based on the input, enhancing the ability to filter noise between events. The second is a scale-invariant event memorization technique, which dynamically determines how an event should be memorized in the hidden state based on its information density, rather than its duration.
3.2.1. Fully Video-aware History Modeling
In existing designs, once trained, the model parameters cannot be flexibly adjusted based on the input frame. Intuitively, if some frame is repetitive or does not contain any events, the design should discard this frame without updating the hidden state, and simply copy the result from the last output. Therefore, we aim to improve all matrix parameters into input-dependent ones. Our design incorporates dynamic adaptability, similar to the attention mechanism in transformers. The attention mechanism adapts to different inputs by adjusting its focus on various parts of the input sequence.
Many efforts have explored that, to model the long-range history, the matrix $A$ should be designed carefully. So, in our design, we only improve matrics $B$ and $C$ to input-dependent ones. Then we detail how to improve the matrix $A$ for long-range history modeling.
We improve the $B$ and $C$ as functions of the input, i.e., $B_{t}=f_{B}(x_{t})$ and $C_{t}=f_{C}(x_{t})$. Consequently, during training, we no longer learn the $B$ and $C$ matrices directly but instead learn how to dynamically generate these matrices based on the input. The hidden state is maintained dependent on the input $x$ to only keep the event features. In our method, we utilize the two linear layers as the function $f_{B}$ and the function $f_{C}$, respectively.
Matrix $A$ controls how information from previous states is transferred to the new state. The design of matrix $A$ determines whether the method retains only recent states (local dependency) or captures all past states. Recently, many efforts have been put into how to improve the ability to memorize long-range history by designing matrix $A$.
Inspired by Work (gu2023mamba, ), we redesign the transform matrix $A$ and its initialization method. First, we enhance matrix $A$ by making it a diagonal matrix, which has been proven to significantly improve the capability to model long-range history. Various efforts have explored alternative initialization techniques for matrix $A$ during training to boost model performance. Following this work, we employ random initialization techniques, which achieve similar results. Therefore, in our C-VUE, we utilize a randomly initialized diagonal matrix $A$ to enable long-range memorization.
3.2.2. Scale-invariance Events Memorization
As shown in Figure 2, different events contain varying numbers of frames, but they may convey a similar amount of information. We aim to use hidden states of similar sizes to memorize events with the same amount of information, even though these events have different time scales.
Inspired by Work (nguyen2022s4nd, ), we can leverage discretization techniques to dynamically resample video frames. The input frames $x_{0}$, $x_{1}$, …, $x_{t}$ can be viewed as uniformly sampled from an underlying signal $f(x)$ with a step size $\Delta$, i.e., $x_{t}=f(t\Delta)$. The standard RNN utilizes matrics $A$ and $B$ to update the state in keeping with $\Delta$ (Equation 1 and 2).
To achieve our design goal, we no longer use video frames as the basic unit for sampling but instead use the events as the basic unit for sampling via a dynamic $\Delta$. Then, we apply the dynamic $\Delta$ on the state update procedure to achieve the dynamic resampling. Note that the step size $\Delta$ is not a real sampling rate, and we do not perform the resampling technique on the video frames. All input frames are processed during the continuous video understanding procedure.
We utilize the discretization technique, i.e., zero-order hold (ZOH), utilized in (gu2023mamba, ) to achieve the above dynamic resampling. Specifically, $A$ and $B_{t}$ can be improved with an input-dependent parameter $\Delta_{t}=f_{\Delta}(x_{t})$:
(5) $\overline{A_{t}}=\exp(\Delta_{t}A)$
(6) $\overline{B_{t}}=(\Delta_{t}A)^{-1}\exp(\Delta_{t}A-I)\cdot\Delta_{t}B_{t}$
$\Delta_{t}$ is calculated based on the input $x_{t}$, thus achieves the awareness of the current event. We use a linear layer as the $f_{\Delta}$ function.
3.3. Temporal Relation-guided Spatial Redundancy Reduction
Videos contain both spatial (within-frame) and temporal (between-frame) information, both critical for accurate understanding. However, as shown in Figure 5, video data contains a vast amount of redundant information in spatial space. Directly performing history modeling introduces lots of redundant computations, leading to poor system performance.
Eliminating the redundancy is challenging. Because, firstly, the event feature is sparse. The event feature in a video sample may be captured by only a small portion of the pixel content. Secondly, object movement causes poor locality at the pixel level. It is hard to accurately understand events merely by analyzing two adjacent regions (by setting thresholds). Therefore, if we perform video understanding at the pixel level like traditional approaches, e.g., MoviNet (kondratyuk2021movinets, ) and Mvitv2 (li2022mvitv2, ), there are lots of redundant computation (as shown in the left part of Figure 6).
Figure 5. Video frames of the balance beam sports, selected from the FineGym99 dataset (shao2020finegym, ). The entire gymnastic movement was captured in only 17 frames, lasting less than 700 milliseconds.
Figure 6. Different ideas to model the video information. Traditional methods associate any two pixels or patches with complexity $\mathcal{O}(n^{2})$ or $\mathcal{O}(m^{2})$, where $n$ and $m$ are the pixel and patch counts. By reducing spatial redundancy, we model video content more efficiently.
Given the issues mentioned earlier, many pixels, such as background pixels in Figure 5, do not contain event information, leading to computational redundancy. Therefore, such a method inevitably leads to lots of computation redundancy. Additionally, due to low inter-frame pixel locality caused by object movement, it is challenging to reduce the quadratic complexity simply by searching nearby pixels.
Traditional video processing methods have carried over into the era of LLM-based video foundation models (damonlpsg2024videollama2, ). The vision encoder break a frame into patches and then correlate these patches across frames in an all-to-all manner. The patch number is $m$, and the complexity is $\mathcal{O}(m^{2})$. This approach makes it difficult for video foundation models to balance accuracy and system performance. Correlating a large number of frames is costly in both the training and inference phases. Specifically, the high computational complexity often requires excessive downsampling, which reduces accuracy. We analyze this issue in Section 6 and show how our design can improve video foundation models.
Our approach introduces a temporal relation-guided spatial redundancy reduction. The key insight is that we can learn the rule to reduce the redundant spatial information, that is not related to the event, by correlating these features in the temporal dimension. For example, understanding a jumping action, as shown in Figure 5, is primarily about observing how the person’s pose and position change between frames, which is unrelated to what the person is wearing. Hence, we can first attempt to capture the specific pose and position information of the objects at each moment, then associate them over time for further understanding.
However, it is challenging, as the types of events in the real world are diverse. We cannot merely retain an object’s pose and position. For instance, in Figure 6, it is crucial to also keep the feature of balance beam present in the environment. In this way, we can accurately comprehend that this is a gymnastics action, rather than a simple jump.
To achieve the reduction goal, we first use a pre-trained powerful feature extractor to extract general-purpose spatial feature, and then use a lightweight projection layer to perform the reduction. Specifically, we utilize the CLIP (radford2021learning, ) to perform the preliminary spatial feature extraction. CLIP is one of the most powerful vision encoders, which serves as the vision encoder in mainstream video foundation models (awais2023foundational, ; madan2024foundation, ).
Then we utilize a linear layer followed by an activation function as the projection layer to reduce the redundancy all-purpose spatial features. The linear layer is trained together with our history modeling layer to reduce the redundancy with the guidance of the temporal relationship. After the reduction, the spatial feature can be used to enable efficient long-range history modeling, as shown on the far right of Figure 6. Thus, the complexity of modeling the relationship between two frames is reduced to $\mathcal{O}(1)$.
3.4. Parallel Training with Frame-weighted Loss
Given the recurrent form of our C-VUE may significantly affect the system performance during the training phase. We adopt a parallel training structure to efficiently train long video samples. Additionally, we design a frame-weighted loss to calculate the loss for each frame, thereby accomplishing prompt video understanding.
3.4.1. Efficient Parallel Training
To implement the redundancy reduction design described in Section 3.3, we start by using a pre-trained vision encoder ($\theta_{\texttt{spatio\_encoding}}$) to extract the general-purpose spatial feature. It is important to note that this vision encoder remains frozen during training. Before performing history modeling, we project the frame embeddings using a linear layer with ReLU activation ($\theta_{\texttt{reduction}}$). Recall that this projection is used to perform spatial redundancy reduction based on the temporal relationship. The $\theta_{\texttt{history\_modeling}}$ is the design proposed in Section 3.2 for adaptive history modeling. We unfold the recurrent state transition process into a convolutional form for parallel computation. Finally, we employ another linear layer with ReLU activation ($\theta_{\texttt{post\_proj}}$) to map the output of each time step to the space of the downstream task. The overall structure of our model is illustrated on the right side of Figure 7.
The training process is detailed in Algorithm 1. Although our system (C-VUE) employs a recurrent design, processing one frame at a time. We achieve high parallelism during training. Specifically, in the vision encoder, as well as during spatial feature reduction and post processing, we treat the temporal dimension as a batch dimension for parallel computation.
Data: Video frames and corresponding labels: $\texttt{X}_{1:n}$ and $\texttt{Y}_{1:n}$
Result: Trained parameters: $\theta$
1 begin
2
3      while not convergent do
             $\texttt{Emb}_{1:n}\leftarrow\theta_{\texttt{spatio\_encoding}}^{\star}(\texttt{X}_{1:n})$ // parallel
             $\texttt{Tok}_{1:n}\leftarrow\theta_{\texttt{reduction}}(\texttt{Emb}_{1:n})$ // parallel
             $\texttt{Ir}_{1:n}\leftarrow\theta_{\texttt{history\_modeling}}(\texttt{Tok}_{1:n})$ // in convolution mode
             $\hat{\texttt{Y}}_{1:n}\leftarrow\theta_{\texttt{post\_proj}}(\texttt{Ir}_{1:n})$ // parallel
4             $\texttt{L}\leftarrow\mathcal{L}(\texttt{Y}_{1:n},\hat{\texttt{Y}}_{1:n})$
5             $\theta\leftarrow\theta-\nabla_{\theta}\texttt{L}$
6
7      return $\theta$
Algorithm 1 The parallel training structure of C-VUE. $\theta_{\texttt{spatio\_encoding}}^{\star}$ is frozen during the training.
Figure 7. The model architecture of our C-VUE.
3.4.2. Frame-weighted Loss for Continuous Understanding
Here we introduce the loss function $\mathcal{L}$ in Algorithm 1. Video datasets are often categorized as either motion-focused or scene-focused ones. The former requires understanding the process of action transitions to interpret the video accurately. For instance, in gymnastics sports datasets (shao2020finegym, ), the initial frames of turning once and twice may look similar, but the later frames are crucial for determining the action category. In contrast, scene-focused tasks can often be distinguished by certain static appearances, such as drinking water or reading (sigurdsson2016hollywood, ).
Video datasets are usually annotated in a temporal-grounded manner. To account for the different characteristics of the video dataset, we propose a frame-weighted loss. The key principle is to compute the loss by taking a weighted average across all frames. Formally, this can be expressed as:
(7) $\mathcal{L}_{v}=W^{T}\times[\mathcal{L}_{f}^{1},\mathcal{L}_{f}^{2},\ldots,\mathcal{L}_{f}^{N}],$
where $W$ is the weight vector of the loss from Frame 1 to Frame $N$.
Users can design different weight vectors based on their datasets. For motion-focused datasets, we recommend a linear growth strategy, where the weight of the loss for the initial frame is relatively low and gradually increases over time within the video frames. Specifically, $W=[\frac{i}{N^{2}}]_{i=1}^{N}$. Conversely, for scene-focused datasets, a uniformly weighted approach is recommended. That is, $W=[\frac{1}{N}]_{i=1}^{N}$.
4. Implementation
Network Architecture. We implement the $\theta_{\texttt{spatio\_encoding}}^{\star}$ with the powerful vision encoder, i.e., CLIP (clip_impl, ). $\theta_{\texttt{temporal\_modeling}}$ is implemented with a single layer Mamba (gu2023mamba, ). Here we utilize Mamba implementation for acceleration. Mamba proposes a selective scanning algorithm, which assumes the order in which it does operations does not matter through the associate property. It calculates the sequences in parts and iteratively combines them to calculate the history modeling in a convolution mode.
Hyperparameter Setting. During the training, we freeze the vision encoder and only train the two projection layers and the Mamba layer. The AdamW (loshchilov2017decoupled, ) is adopted as the optimizer. The learning rate changes dynamically, using Cosine annealing with warm restarts. This method adjusts the learning rate between 1e-3 and 1e-6 every 10 epochs. Aside from these adjustments, the other hyperparameters are kept at their default settings as provided in PyTorch v2.2.0. Regarding the training process of the network, it requires approximately 100 to 200 epochs to converge in different datasets. We train these models on a server with 8 RTX4090 GPU cards, and it takes about 3 to 6 hours to train a model on different datasets.
Optimization Deployment. To ensure real-time operation and cross-platform deployment, we have utilized compilation and quantization techniques on our model. We have encoded the model file into $C$ code to guarantee cross-platform deployment, and ultimately achieved deployment on a MacBook Air (M3, 2024), Jetson AGX Orin 64 GB, and a PC with an RTX 4090 GPU card.
We implemented our vision encoder CLIP (radford2021learning, ) based on the llama.cpp framework (llama_cpp_impl, ), completed the compilation, and carried out INT8 quantization. The existing CLIP.cpp (clip_cpp_impl, ) framework based on llama.cpp for compiling CLIP only supports CPU backend, hence it can’t utilize the GPU on mobile devices. Other mainstream frameworks lack good compatibility for cross-platform deployment. For example, OpenClip (open_clip_impl, ) and ONNXRuntime-GPU (onnx_impl, ) lack support for using the CLIP model with Jetson GPU. We have expanded support for ggml in our implementation, enabling GPU-based inference on all the aforementioned platforms, which resulted in significant system performance improvement. Please refer to Figure 9. The official implementation of Mamba (mamba_impl, ) only supports the PyTorch framework with the CUDA backend. We have compiled our Mamba part into a $C$ program, thereby allowing its deployment on various platforms.
5. Evaluation
5.1. Experiment Setups
Evaluation Goals. We aim to evaluate five capabilities of the system:
Capability 1: Understanding of events with various lengths.
Capability 2: Prompt response as more frames are received.
Capability 3: Generality for long-tail data.
Capability 4: Understanding of continuous input events.
Capability 5: Long-range memorization.
Figure 8. Datasets used in experiments. S represents a sample; a line segment represents the duration of an event; E represents a label. In Charades and Multi-label FineGym, a frame may have multiple labels.
Datasets. We utilize five datasets to evaluate different capabilities. FineGym99 (shao2020finegym, ) (For capabilities 1 & 2). It is a gymnasium video dataset. Each video clip is labeled with a gymnastics movement. The duration of events varies from 2 to 50 seconds.
FineGym288 (shao2020finegym, ) (For capabilities 1, 2, & 3). It is an extended version of FineGym99, which enhances its by adding categories. FineGym288 has the natural long-tail distribution.
Charades (sigurdsson2016hollywood, ) (For capabilities 1, 2, & 5). It is a multi-label dataset, composed of daily indoor activities with an average length of 30 seconds. Each video clip consists of about 800 frames on average, varying from 32 to 3,944 frames. We have re-annotated the video samples for continuous video understanding as Figure 8(b).
All-event-in-one FineGym (For capabilities 1, 2, & 4). It is constructed based on the FineGym99 as shown in Figure 8(c). We concatenated all the video clips from FineGym99’s testing dataset, forming a single long video lasting 3.5 hours and consisting of over 440k frames.
Multi-Label FineGym@k (For capabilities 1, 2, & 5). It is constructed based on the FineGym99 as shown in Figure 8(d). The new video clips are constructed by randomly concatenating $k$ events of FineGym99 together and re-annotating them.
Baseline Methods. We utilize five baselines to execute continuous perception, whereby the method predicts an outcome at each timestep. Baselines 1, 2, and 3 are constructed following different video processing paradigms. Baselines 4 and 5 are state-of-the-art (SOTA) methods for balancing accuracy and efficiency for video classification.
Baseline 1: Transformer (vaswani2017attention, ). The overall architecture mirrors ours for a fair comparison, except that we replace our history modeling layer with a transformer of equivalent parameter size.
Baseline 2: ConvPooling (kar2017adascan, ). The overall architecture mirrors ours for a fair comparison, except that we replace the history modeling layer with a convPooling of equivalent parameter size. The convPooling consists of a 1D convolution and a 1D adaptive max pooling.
Baseline 3: LSTM (hochreiter1997long, ). This design keeps the overall structure consistent with ours for a fair comparison, substituting the history modeling layer with an LSTM of equal parameter magnitude.
Baseline 4: MoviNetA0 (kondratyuk2021movinets, ). This is a Mobile Video Network specifically designed for efficient video recognition. This method applies down-sampling when processing videos. Drawing from the experience of using this model on the Charades dataset, we similarly down-sample the video to 6 fps for video preprocessing.
Baseline 5: Mvitv2(li2022mvitv2, ) This is an improved multiscale vision transformer for both classification and detection tasks. This method necessitates the sampling of any given video to a fixed number of frames for video classification tasks. We trained the baseline that takes as input 16 frames at a time, which is one of the default settings.
Input Strategy. We utilize three typical input strategies to perform the evaluation.
S1: All frames feeding. All frames from frame 1 up to frame $N$ are fed into the baseline method to provide the necessary context for achieving continuous perception at timestep $N$.
S2: Sliding windowed feeding. All frames between the $(N-m)$-th frame to the $N$-th frame are fed into the baseline at timestep $N$, where $m$ denotes the sliding window size.
S3: Single frame feeding. Only frame $N$ is supplied to the baseline method.
Metrics. In this work, we utilize various metrics, namely Accuracy, Mean Average Precision (mAP), and the Jaccard Index for single-label and multi-label video classification tasks. Furthermore, we introduce a novel metric referred to as the Early Detection Rate (EDR), designed to evaluate the model’s ability to detect events promptly within videos. The definition is $\mathtt{EDR}={(E_{pred}-E_{gt})}/{E_{N}}$, where $E_{pred}$ refers to the frame index at which the label is detected. $E_{gt}$ represents the ground truth frame index, that is, the index of the frame where the event begins. $E_{N}$ signifies the total number of frames in the event.
Table 1. The accuracy and EDR results of different baselines. The #Param counts the parameters that need to be trained. Baselines 1, 2, 3, and our method have the same pre-trained clip with a parameter size of 149M.
FineGym99 FineGym288 Charades
Methods #Param Acc (%) EDR (%) Acc (%) EDR (%) $\text{mAP}_{\text{all}}$ (%) $\text{mAP}_{\text{last}}$ (%) EDR (%)
Baseline 1 (Transformer) 210.5K 82.3 48.20 76.6 53.20 17.2 21.4 52
Baseline 2 (ConvPooling) 367.3K 82.6 38.40 75.4 41.40 15.8 21 43.3
Baseline 3 (LSTM) 218K 84.1 23.40 78.4 25.80 15.33 17.8 84
Baseline 4 (MoviNet) 2.1M 19.7 41.60 13.9 44.20 3.9 5.8 32
Baseline 5 (Mvitv2) 34.3M 6.4 20.50 3.1 24.10 5.3 7.3 99.9
Ours 202K 86.1 22.50 81.3 25.10 18.4 22.1 56.3
5.2. Evaluation of All Five Capabilities
In Table 1, we present the outcomes from several baseline models across different datasets. We first introduce each column’s meaning and statistical methods, followed by an evaluation of our methodology and findings.
The #Param column signifies the size of a model’s parameters. Owing to the varying category numbers in the three datasets, there is a marginal difference in the parameter count in the final layer for different tasks, approximately a 10k difference in parameters. Here, we reference the parameter count when classifying the FineGym99 dataset. Baselines 1, 2, and 3, as well as our model, all utilize the same CLIP model with a parameter size of 149M. When excluding the CLIP, our design exhibits the smallest parameter count.
We chose the most natural input strategy for each baseline to assess the accuracy. For batch-based methods, i.e., Baselines 1, 2, 4, and 5, we adopted the all-frame feeding strategy. Although this significantly increases computational overhead, it enables these methods to have sufficient context to understand the video content. Both Baseline 3 and our method utilize a streaming processing paradigm, for which we selected the single-frame feeding strategy.
In the evaluation of single-label classification datasets, we utilize Accuracy (Acc) as the primary measure of correctness. This accuracy is ascertained by verifying whether the correct action is detected within a given video. The EDR is calculated as introduced in Section 5.1(Metrics). When dealing with multi-label classification datasets, two variants of mean Average Precision (mAP) are employed to gauge correctness. The $\text{mAP}_{\text{all}}$ metric is determined by considering predictions of all frames in the video. The $\text{mAP}_{\text{last}}$ entails calculations based solely on the prediction results of the final frame in the video. It’s important to note that the label of the last frame includes labels for all actions present within the current video. Each label’s EDR is computed separately and then averaged. If a method fails to detect a particular label, this failure is taken into account during computation by adding a 100% increase to the result for that category. This signifies that the correct label was not detected even upon the conclusion of the video.
Below we present our experimental analyses and the subsequent conclusions:
Superior Accuracy: Our method consistently outperforms all baseline methods in terms of accuracy across all tasks, improving metrics by ranges of 2 to 79.7, 2.9 to 78.2, 1.2 to 14.5, and 0.7 to 16.3. The leading accuracy demonstrates that our C-VUE can effectively understand events of different scales (Capability 1). The superior accuracy on the FineGym288 dataset underscores our method’s generalization ability for long-tailed data (Capability 3)).
Prompt Event Detection: Our approach exhibits an outstanding ability to promptly detect events (Capability 2) and continuously processing capability (Capability 4) by EDR. For single-label tasks, events are detected within the first quarter of their occurrence. For complex multi-label tasks, all events are identified halfway through the video.
Memorization Capability: Within the multi-label dataset Charades, the $\text{mAP}_{\text{last}}$ denotes the ability to recognize all events in the video by its conclusion. Notably, Baselines 1 and 2 attained such accuracy by taking as input all video frames simultaneously. Our method achieves a higher accuracy based on the last frame and implicit memorization of past frames. Notably, our approach achieves 1.24 times the accuracy of Baseline 3, which employs LSTM for continuous processing. Results show that our C-VUE shows a better long-range memorization ability (Capability 5).
Optimal Use of Trainable Parameters: Notably, our method retains superior performance while utilizing a minimal number of training parameters, amounting to only 202K. This represents a drastic reduction when compared to the parameter requirements of baseline models, with differences ranging from 8.5K to 34.1M.
Robust Stability Across Varied Tasks: Our method showcases robust stability in tackling diverse tasks. It performs uniformly well across differing task scales (including those with varied event durations), varied video characteristics (motion-focused and scene-focused video), and differing classification objectives (such as single-label and multi-label classifications). No baseline methods are currently able to match this level of comprehensive performance.
Considering the inadequate accuracy demonstrated by Baselines 4 and 5, subsequent experimental evaluations will only involve Baselines 1, 2, and 3, along with ours.
Table 2. Evaluation results on All-event-in-one FineGym. (The parenthesized number of S2 signifies the window size, while 725 refers to the maximum frame count of a single event.)
Method Baseline 1 Baseline 3
Input
Strategy
S3
S2
S2
S2
S1 S3
Acc (%) 4.1 6.6 4.2 6.2 - 7
Method Baseline 2 Ours
Input
Strategy
S3
S2
S2
S2
S1 S3
Acc (%) 4.7 11.8 13.4 11.3 7 14.8
5.3. Further Evaluation of Capabilities 1 & 4
We assess the continuous perception capability with the All-event-in-one FineGym dataset, a 3.5-hour video described in Section 5.1. All baseline methods use the model trained on the FineGym99 dataset, whose accuracy evaluated on a single video clip is reported in Table 2. These methods are evaluated with different input strategies. For batch processing approaches (Baselines 1 and 2), we test all three strategies. Regarding streaming processing methods (our C-VUE and Baseline 3), we use the single frame feeding strategy (Strategy 3). Both our C-VUE and Baseline 3 benefit from the model’s memorization of past video frames. During continuous perception, each frame is processed only once without redundant input. The Acc metric is calculated by considering predictions of all frames in the video.
Baseline 1’s transformer structure results in slower processing as the input sequence grows. Therefore, Baseline 1 faced challenges during evaluation with input strategy 1 due to increasing input sequence length over time. Detailed system performance is provided in Figure 4.
Based on the results, we draw the following two conclusions. First, our C-VUE demonstrates the best continuous perception ability (Capability 4). It surpasses batch processing methods ranging from 0.5 to 10.7. Its accuracy is twice that of traditional LSTM-based baseline methods. Second, our approach can effectively adapt to events of varying scales (Capability 1). In the hours-long video, events lasted from tens of frames to hundreds of frames. Our C-VUE achieves the best accuracy, which outperforms the current sliding-window and state-assist methods.
5.4. Further Evaluation of Capability 5
We further profile the memorization abilities on three multi-label Finegym datasets to evaluate our memorization capability. The k value represents how many events are concatenated to form a testing data sample. Therefore, a higher k value indicates a longer video with more labels.
All baseline methods use the model trained on the FineGym99 dataset. Baselines 1 and 2 are evaluated with input strategy 1, where they can achieve the best accuracy. We utilize three metrics. The Jaccard Index [^1] The threshold ‘th’ signifies that in the model’s output vector, values greater than ‘th’ are considered as detecting a particular category, otherwise not recognized. is calculated as the intersection of predicted results and ground truth divided by their union. We report the results in Table 3 and draw the following conclusions.
Our C-VUE demonstrates an superior memorization capabilities (Capability 5). Across all configurations and metrics, our method surpasses the baseline approaches. Our C-VUE excels at tasks involving frame-by-frame understanding. Surprisingly, baseline methods that process frames in batches (Baselines 1 and 2), incorporating all frames at each time step, exhibit notably lower accuracy compared to streaming methods (Ours and Baseline 3). This difference arises from the fact that streaming methods prioritize the present frame while retaining information from past frames in streaming data processing.
Table 3. Evaluation results on three Multi-label FineGym@k datasets.
Metrics Methods k=1 k=3 k=5
Jaccard Index (th=0.5) Baseline 1 0.0351 0.0172 0.0087
Baseline 2 0.0194 0.0281 0.0313
Baseline 3 0.0906 0.0484 0.0287
Ours 0.1210 0.0578 0.0341
Jaccard Index (th=0.3) Baseline 1 0.0803 0.04396 0.0228
Baseline 2 0.0101 0.0152 0.0200
Baseline 3 0.1426 0.0880 0.0554
Ours 0.1752 0.0957 0.0604
Jaccard Index (th=0.1) Baseline 1 0.1324 0.0802 0.0441
Baseline 2 0.0101 0.0152 0.0200
Baseline 3 0.1428 0.1254 0.0935
Ours 0.1846 0.1334 0.0936
mAP Baseline 1 0.1590 0.1196 0.0905
Baseline 2 0.0465 0.0374 0.0353
Baseline 3 0.2338 0.1645 0.1277
Ours 0.2678 0.1739 0.1379
ranking AP Baseline 1 0.3879 0.3296 0.2771
Baseline 2 0.4533 0.4338 0.4108
Baseline 3 0.5856 0.4906 0.4250
Ours 0.5894 0.4916 0.4346
Figure 9. Processing speed of understanding a video with 10k frames on different platforms.
5.5. Real-time Inference on Edge
We measure the system performance on three mobile platforms with two long video clips. The long video clips were excerpted from the All-event-in-one FineGym dataset, one containing 10k frames and the other 100k frames. We conducted each experiment three times and calculated the average. In Figure 9, we present the system performance of different baselines processing 10k frame videos on the three platforms. Additionally, in Figure 4, we report the system performance of different baselines processing 100k frame videos on a PC equipped with an RTX4090. FPS is calculated by dividing the total number of video frames by the total processing time. Our experimental findings are as follows.
Understanding the video stream in real time. In Figure 9, it can be observed that our method achieves real-time processing (¿30fps) on different platforms. Even on Macbook Air and Jetson, our speed is 2 to 20 times faster than the baseline methods. This improvement stems from two aspects. Firstly, our streaming processing method eliminates the need to use the model through all feeding or sliding window approaches. Secondly, in terms of system implementation and optimization, recall that mainstream frameworks, e.g., OpenClip, clip.cpp, and ONNXRuntime-GPU, lack support for using the CLIP model with Jetson GPU. Therefore, baselines can only utilize the CPU for computation. Our prototype can enable GPU computation on both Macbook Air and Jetson.
No significant performance degradation as the frame number increases. As shown in Figure 4, with an increase in input sequences, the performance of Baselines 1 and 2, using the all frame feeding and sliding windowed feeding strategy, significantly declined. When processing a 100k frame video, Baseline 1 drops to a speed of 9fps.
6. Case Study: Video Encoder for LLM-based Video Foundation Models
Figure 10. Applying our history modeling into VideoLLaMa2 as a video encoder.
In this section, we discuss how the design of our system, C-VUE, improves video foundation models. As a case study, we analyze one of the SOTA video foundation models, VideoLLaMa2 (damonlpsg2024videollama2, ) (left side of Figure 10). During video processing, VideoLLaMa2 first samples $N$ frames from the video. These frames are then converted into $M$ tokens by a multimodal projector (denoted as ”Proj.” in the figure). Simultaneously, the user’s text prompt is tokenized, combined with the video tokens, and input into the LLM for inference.
6.1. Problem Analysis
VideoLLaMa2 introduces an STC projector as the projection layer. As shown in the middle part of Figure 10, the STC projector utilizes a batch-based idea to understand the video without reducing the redundancy. Specifically, it performs 3D convolutions on the feature of $N$ sampled frames. The design of the STC projector limits VideoLLaMa2’s effectiveness and performance. Therefore, it struggles to support higher sampling rates because 3D convolutions on a larger batch demand significant computational resources. As a result, only two versions of the model are available: one that processes 8 frames and another that processes 16 frames. Regardless of the video length or event scale, this fixed frame sampling inevitably reduces the model’s accuracy.
6.2. Our design
We adopt the design of our C-VUE to build a new projection layer, named the C-VUE projector, which replaces the STC projector. This new projection utilizes the long-range history modeling to enable a high sampling rate in a remarkably efficient manner. Figure 10 (right) illustrates our design.
When a video frame is input, the projector first processes the spatial information within the frame by applying mean pooling to patch embeddings. Next, the history modeling layer processes the history information, capturing relationships over time. Its processing speed remains consistent regardless of video length, making it more efficient.
Table 4. Evaluation results on the VideoLLaMa2.
VideoChatGPT
Corr. (5-point)
ActivityNet
Acc (%)
MSVD
Acc (%)
VideoLLaMa2 1.95 38.46 44.6
Ours 2.41 39.65 48.6
6.3. Experiments
Implementation. Using the official videollama2 codebase (video_llama_impl, ), we implemented our C-VUE projector. This projector consists of mean pooling, a linear layer with ReLU, and the history modeling layer of our C-VUE with a code dimension of 2048, followed by another linear layer with ReLU. The total parameter count of our projector is 450 M. For comparison, the STC projector with 8 frames has 1.8 B parameters, which is four times larger.
We fine-tuned and evaluated the model following the official guidelines. All hyperparameters were set to the official default values. During fine-tuning, only the large language model (LLM) and the projection layer were trained. We used the videoChatGPT dataset (maaz2023video, ) for fine-tuning both the original model structure and our proposed model structure. All experiments were performed on a machine with eight A600 GPUs. Due to limited training resources, we used the 8-frame model as the baseline, while our method extended the training sample length to 128 frames.
Results. Testing was conducted on the videoChatGPT dataset (in-distribution), ActivityNet-QA (yu2019activityqa, ) (zero-shot), and MSVD-QA (xu2017video, ) (zero-shot), and the results are reported in Table 4. For the VideoChatGPT dataset, correctness was measured on a 5-point scale. Our method outperformed the baseline by 0.46 points. On the other two zero-shot datasets, our approach achieved improvements of 1.19% and 4%, respectively. These results demonstrate that our projector design effectively improves model accuracy, highlighting a new design concept for video projection for video foundation models.
7. Related Work
Video Analytics & Understanding. Video analytics and understanding involve the automatic extraction of valuable information from video including classification (wu2020emo, ), detection (tong2021large, ; jiang2021flexible, ), tracking (voigtlaender2019mots, ), identification (yi2020eagleeye, ), and privacy protection (wu2021pecam, ). Many approaches treat continuous video analysis as a task of image understanding on video frames. Adaptive frame sampling (hwang2022cova, ; moll2022exsample, ), cooperation between small and large models (liu2020continuous, ), and collaboration across multiple computation resources (chen2015glimpse, ; liu2019edge, ; damonlpsg2024videollama2, ) are proposed to enable efficient video analytics on video frames.
To analyze the events that involve multiple frames, 3D convolution is used to perform the video understanding (slowfast2019fei, ). However, this approach can cause significant system performance issues (zhu2020comprehensive, ). To improve efficiency, MoviNet (kondratyuk2021movinets, ) creates efficient 3D convolution architectures. Many efforts also explore temporal modeling without 3D convolution (Lin_2019_ICCV, ; Jiang_2019_ICCV, ; Li_2020_CVPR, ). MViTv2 (li2022mvitv2, ) uses improved pooling attention for spatiotemporal recognition to improve the attention’s efficiency. These methods improve video understanding by focusing on relationships in a short video clip. They still struggle with system performance issues in long-form and streaming videos.
Video Foundation Models. Recent studies have explored building video foundation models using large language models (LLMs). These models achieve impressive results in multiple video understanding tasks (madan2024foundation, ).
Methods like VideoLaVIT (jin2024video, ), VideoLLaVA (lin2023video, ), Video-LLaMA (zhang2023video, ), and Video-LLaMA2 (damonlpsg2024videollama2, ) use multi-modal pre-training to align visual information with LLMs through projection layers. VideoChatGPT (maaz2023video, ) improves video-based conversations by integrating visual encoders with LLMs and instruction tuning. The main challenge in video understanding is handling memory requirements and model capacity, especially for long videos. LaViLa (zhao2023learning, ) addresses this by turning LLMs into narrators, allowing them to generate detailed descriptions of long videos from visual inputs. ChatVideo (wang2023chatvideo, ) and MM-VID (lin2023mm, ) convert videos into text for better comprehension. However, these methods still struggle to model long-range temporal information in streaming videos.
Alternatives to Foundation Models for Video Understanding. The series of works on Mamba have gained significant attention for overcoming the high computational complexity of Transformers as LLMs. Recently, Mamba has been applied to various computer vision tasks, showing early progress (liu2024vision, ). However, its application is still in its infancy. Mamba-ND (li2024mambandselectivestatespace, ) is the first to demonstrate Mamba can handle high-dimensional data and experimentally show the effectiveness in video action recognition. VideoMamba (li2024videomambastatespacemodel, ) builds an SSM-based model for video understanding. It shares a similar design with VideoLlama (damonlpsg2024videollama2, ) but replaces the Transformer-based LLM with a Mamba-based one for efficiency purposes. VideoMamba Suite (chen2024video, ) further evaluates Mamba’s performance by integrating it into existing video understanding architectures, showing its potential to handle complex spatial-temporal dynamics.
Unlike these works that use “all frames at once” or “fixed window size” strategies for Mamba and Transformer video understanding, we propose a lightweight continuous video understanding design with each frame as the input to the state-assistance module.
8. Discussion & Conclusion
Discussion. Our work presents a solution for continuous video understanding, but there are still unexplored issues we hope to address with the community in the future. The memorization upper bound of the state modeling approach has not been fully explored. Although several improvements were made, a fixed-size hidden state will fail to model infinitely long videos. Future work should figure out the modeling capacity of such methods. Additionally, the generalization ability of these approaches to more and more tasks remains unclear, which is crucial for designing new video foundation models based on this design.
Conclusion. Continuous video understanding is important for complex multimodal human-computer interaction. However, achieving this on resource-limited edge devices is challenging. This paper proposes C-VUE, an adaptive state modeling approach, for continuous video understanding. Experimental results show that C-VUE can achieve real-time continuous video understanding on typical edge platforms. As a case study, we apply C-VUE’s design to improve video foundation models and obtain a large accuracy improvement.

### v2

@@
-Making Every Frame Matter: Continuous Video Understanding for Large Models via Adaptive State Modeling
-Abstract.
-Video understanding has become increasingly important with the rise of multi-modality applications.
-Understanding continuous video poses considerable challenges due to the fast expansion of streaming video, which contains multi-scale and untrimmed events.
-We introduce a novel system, C-VUE, to overcome these issues through adaptive state modeling.
-C-VUE has three key designs.
-The first is a long-range history modeling technique that uses a video-aware approach to retain historical video information.
-The second is a spatial redundancy reduction technique, which enhances the efficiency of history modeling based on temporal relations.
-The third is a parallel training structure that incorporates the frame-weighted loss to understand multi-scale events in long videos.
-Our C-VUE offers high accuracy and efficiency.
-It runs at speeds ¿30 FPS on typical edge devices and outperforms all baselines in accuracy.
-Moreover, applying C-VUE to a video foundation model as a video encoder in our case study resulted in a 0.46-point enhancement (on a 5-point scale) on the in-distribution dataset, and an improvement ranging from 1.19% to 4% on zero-shot datasets.
-1. Introduction
-Video understanding (madan2024foundation, ; chen2024video, ; huang2018makes, ; zolfaghari2018eco, ; wu2021towards, ; tang2023video, ) is an important and long-standing application for edge and mobile computing.
-It requires capturing spatio-temporal features from an input video to conduct high-level tasks, such as action localization, activity recognition, video question answering, and video captioning.
-These tasks are widely used in edge scenarios, including robotics, augmented reality, embedded devices (e.g., Kinect and Wii remote), and IoT devices.
-Due to user data privacy, responsiveness requirements, and the high cost of video transmission to the cloud, real-world video understanding is typically processed on edge devices.
-Challenges.
-Compared to video processing tasks only relying on frame spatial features (wu2020emo, ; tong2021large, ; jiang2021flexible, ; voigtlaender2019mots, ; yi2020eagleeye, ; wu2021pecam, ), e.g., object detection, segmentation, and tracking, video understanding on edge is much challenging due to the following reasons.
-(i) Video understanding needs to capture the spatio-temporal features of an event on different temporal scales (zhao2018trn, ; shao2020finegym, ).
-Figure 2 shows an indoor activity clipping from Charades dataset (sigurdsson2016hollywood, ).
-The temporal durations of activities vary greatly from around 155 frames (putting a bag) to 553 frames (making a sandwich).
-The challenge lies in modeling this variability effectively: shorter events demand quick, precise attention, while longer activities require retaining relevant context over extended periods, complicating feature extraction across different timescales.
-(ii) Video understanding normally needs to capture meaningful features from the raw streaming videos, which are untrimmed and dominated by noise and redundancy in both spatial and temporal dimensions.
-As seen in Figure 2, the pixels that are related to an activity are very sparsely distributed in the video.
-The understanding may be distracted by the dominant unrelated pixels.
-(iii) Video understanding on edge must guarantee a timely understanding of the scene while managing the resource constraints of typical edge devices.
-For instance, robots or embodied AI (duan2022survey, ) systems must continuously perceive and comprehend dynamic scenarios through video input to carry out reasoning (strafforello2023current, ) and task planning (gadre2022continuous, ).
-Figure 1. C-VUE: continuous video understanding on streaming video. As each frame is received, C-VUE updates its hidden state and outputs the events that have occurred so far.
-Figure 2.
-We utilize the video stream, along with annotations from the Charades dataset (sigurdsson2016hollywood, ), to illustrate our motivation.
-C-VUE aims to understand ”What has happened so far?” at any moment in a streaming video.
-Figure 3.
-Typical designs to process frames of the video stream for continuous video understanding.
-$\texttt{E}_{i}$ signifies an event while $\texttt{f}_{i}$ depicts the frame at time $t_{i}$.
+Making Every Frame Matter: Continuous Activity Recognition in Streaming Video via Adaptive Video Context Modeling
+Abstract
+Video activity recognition has become increasingly important in robots and embodied AI.
+Recognizing continuous video activities poses considerable challenges due to the fast expansion of streaming video, which contains multi-scale and untrimmed activities.
+We introduce a novel system, CARS, to overcome these issues through adaptive video context modeling.
+Adaptive video context modeling refers to selectively maintaining activity-related features in temporal and spatial dimensions.
+CARS has two key designs.
+The first is an activity spatial feature extraction by eliminating irrelevant visual features while maintaining recognition accuracy.
+The second is an activity-aware state update introducing dynamic adaptability to better preserve the video context for multi-scale activity recognition.
+Our CARS runs at speeds $>$30 FPS on typical edge devices and outperforms all baselines by 1.2% to 79.7% in accuracy.
+Moreover, we explore applying CARS to a large video model as a video encoder.
+Experimental results show that our CARS can result in a 0.46-point enhancement (on a 5-point scale) on the in-distribution video activity dataset, and an improvement ranging from 1.19% to 4% on zero-shot video activity datasets.
+1 Introduction
+Video activity recognition [26, 7, 16, 48, 42, 37] is an important and long-standing application.
+These tasks are widely used in advanced human-computer interactions (HCI), including robotics and augmented reality [9, 35, 11].
+To enable advanced HCI, real-time recognition of streaming video on resource-constrained mobile or edge devices is crucial.
+Activity recognition relies heavily on video context, making it difficult to determine actions from a single frame.
+Therefore, a continuous video activity recognition system must maintain video context, allowing it to infer activities in new frames based on the accumulated context.
+However, achieving efficient continuous recognition in streaming video is challenging for the following reasons.
+Figure 1: An example of continuous video activity recognition in streaming video. CARS takes every frame as input and continuously recognizes human activities.
+(i) Diverse Temporal Scales of Activities.
+Continuous video activity recognition needs to capture the spatio-temporal features of activity on multiple temporal scales [46, 30].
+As shown in Figure 1, activities at different scales may have dependencies or causal relationships, such as taking bread from a bag as part of making a sandwich.
+The system must continuously recognize activities of varying lengths based on newly received video frames.
+(ii) Accumulating Redundancy in Streaming Video.
+Continuous video activity recognition normally needs to capture meaningful features from the raw streaming videos, which are dominated by noise and redundancy.
+As illustrated in Figure 5, activity-related pixels are sparsely distributed in the video.
+Without effective online reduction, the accumulation of redundant features can quickly overwhelm the activity-related features as the video progresses.
+(iii) Computational Efficiency Under Resource Constraints.
+Continuous video activity recognition must guarantee a timely recognition of the scene while managing the resource constraints of typical edge devices.
+For instance, robots or embodied AI [9] systems must continuously perceive dynamic scenarios through streaming video input to carry out reasoning [35] and task planning [11].
+Existing video recognition efforts have explored numerous strategies for maintaining video context, yet none have been able to address all the challenges simultaneously.
+Works, using “all frames at once” strategy (Figure 3(a)), perform activity recognition on every frame of the video, which theoretically yields the highest accuracy [10, 19].
+However, its efficiency is the biggest concern.
+Other works use “sliding window” strategy (Figure 3(b)) to reduce computational cost [23, 17, 21, 22].
+However, these methods struggle to accurately recognize activities at different scales.
+Some approaches use a “hidden state assistance” design (Figure 3(c)), where a hidden state compresses video context to enable more efficient activity recognition.
+However, existing state-assisted methods do not effectively eliminate spatiotemporal redundancy in the streaming video, resulting in inefficient utilization of the hidden state and low accuracy.
+Figure 2: Difference between our idea and existing solutions on maintaining the hidden state.
+The all-frame-at-once approach suffers from efficiency degradation over time as data volume grows, while the sliding window method often faces challenges in selecting optimal window sizes or sampling strategies, compromising accuracy.
+Therefore, our design builds on the existing state-assistant framework, which is efficient—processing one frame at a time—and accurate—adaptively maintaining video context through carefully designed hidden states.
+As illustrated in Figure 2, we argue that existing state-assistant works lack activity-aware design in both temporal and spatial dimensions.
+They tend to compress all video features, including the redundant background features, into the hidden state, resulting in inefficient use of hidden states, poor preservation of video context, and low recognition accuracy.
+Our core idea is to selectively compress only activity-related features in temporal and spatial dimensions, which we name adaptive video context modeling.
+Figure 3: Typical designs for streaming video activity recognition.
 The “All Frames at Once” design processes all video frames at every moment.
-The “Sliding Window” design incorporates only the latest $m$ frames.
-The “State Assistance” relies on a single frame at a time and archives previous data through a hidden state.
-Our adaptive state modeling is an improved state assistance design, excelling in long-range history modeling via a video-aware technique.
-Current status of technology.
-Past works explore different DNN methods for video understanding, including 3D CNN (tran2015learning, ), RNN (e.g., LSTM) (hochreiter1997long, ), Transformer (vaswani2017attention, ), and the recent Mamba (gu2023mamba, ).
-However, none of the past works can mitigate all the challenges at the same time.
-To capture the features of various temporal scales, a straightforward method is to input all the past frames into the model at each time step, as illustrated in Figure 3(a).
+The “Sliding Window” design incorporates only the latest (or sampled) $m$ frames.
+The “State Assistance” relies on a single frame at a time and maintains video context through a hidden state.
+In this paper, we propose CARS for continuous video activity recognition in the streaming video, utilizing the adaptive video context modeling, with two key designs.
+First, activity spatial feature extraction (Section 3.1).
+To reduce redundant visual information, we employ a data reduction strategy that operates without requiring supervision from masks or bounding boxes of activity-related features.
+During training, we maximize feature elimination from the video input while preserving recognition accuracy.
+This adversarial learning process allows our CARS to discard spatial features irrelevant to the activity.
+Second, activity-aware state update (Section 3.2).
+Existing methods struggle to effectively maintain video context for multi-scale activities.
+Our design integrates a dynamic adaptability, akin to the attention mechanism in transformers.
+We dynamically recalculate the weight parameters based on the current input frame during the input processing, hidden state update, and output generation.
+We evaluate CARS across various activity recognition tasks, comparing it with widely-used models for video recognition.
+Results show that with significantly fewer trainable parameters (ranging from 8.5k to 34.1M less than the baselines), CARS achieves accuracy improvements of 1.2% to 79.7% over baseline methods.
+Regarding latency, we evaluate CARS on typical edge devices (NVIDIA Jetson Orin, Apple MacBook, and an Intel PC with NVIDIA 4090 GPU).
+It achieves $>$30FPS on all these devices, up to 20$\times$ faster than the available inference systems.
+To further explore the capability of our design, we apply CARS as a video encoder to a large video model to improve its activity recognition capability (Section 5).
+Specifically, we replace the visual encoder of the SOTA large video model VideoLlama2 [8] by CARS.
+The results demonstrate a 0.46-point improvement (on a 5-point scale) on the in-distribution dataset, and an improvement ranging from 1.19% to 4% on zero-shot datasets.
+Contribution. We summarize our contribution as follows:
+  • We propose a continuous video activity recognition system, CARS, based on adaptive video context modeling.
+It promptly recognizes the activities in incoming video frames by efficiently maintaining video context for multi-scale activities.
+  • Our CARS incorporates two key techniques to enhance the utilization efficiency of hidden states, i.e., minimizing redundancy from activity-irrelevant spatial information, and improving adaptability to streaming input while maintaining state.
+  • Experiments show that CARS surpasses all baselines in streaming video activity recognition, using fewer parameters and achieving over 30 FPS on typical edge devices, which can enable on-device video recognition for embodied AI.
+We also showcase that CARS can enhance the large video models by serving as the video encoder.
+2 Background & Related Works
+Various approaches strive to achieve continuous activity recognition, but none have solved this challenging task.
+Figure 4: Processing speed of video activity recognition on RTX4090 (a 100k-frame video).
+Baselines 1(A) and 2(A) represent Transformer and Convpooling-based solutions using the “All Frames at Once” design.
+Baselines 1(F) and 2(F) employ the “Fixed window size” design (the window size is 100).
+Baseline 3 denotes an LSTM-based “State Assistance” design.
+All Frames at Once.
+A straightforward method is to input all the past frames into the model at each time step, as illustrated in Figure 3(a).
 However, it is certainly too costly for model inference and training.
-Take a Transformer model as an example, with the number of streaming frames $n$ grow, the computation cost of video understanding is $O(n^{3})$.
-Even with the many linear alternatives of Transformers, such as Linformer (wang2020linformer, ), RetNet (sun2023retentive, ), Mamba (gu2023mamba, ), and RWKV (peng2023rwkv, ), the cost is still $O(n^{2})$.
-The most commonly used technique to control the cost is to fix the input size of the model even with the growing number of frames, such as sliding window or sampling, shown in Figure 3(b).
-For example, TSN (wang2016temporal, ) first divides the past video into segments and selects a random snippet from each segment for activity recognition.
-VideoLlama2 (damonlpsg2024videollama2, ) only samples 8 or 16 frames as input for any video length.
-This method has to ignore many detailed features.
-There are also works (shi2015convolutional, ; huang2020efficient, ; zhou2018temporal, ), shown in Figure 3(c), that only input one current frame into a model (e.g., CNN) first to extract spatial features, and then conduct feature aggregation (e.g., through pooling (kar2017adascan, ) or LSTM (hochreiter1997long, )) among frames over time to capture temporal dependencies by hidden states.
-This is the most efficient method for inference and training, and all frames can be processed with no one lost.
-This method, however, struggles to capture the temporal relations in long range, as the fundamental limitation discussed in (shao2020finegym, ; wang2016temporal, ).
-Current works’ limitations highlight the difficulty in creating a continuous video understanding method that meets the above three challenges.
-An ideal strategy should keep a fair balance between the accuracy of the Transformer and the efficiency of the LSTM, interpreting multi-scale events in rapidly growing videos.
-1) It needs to correctly keep track of past state when forming decisions.
-2) It must operate in real time on typical edge devices, importantly, without reducing processing speed as video length expands.
-Key techniques.
-To address the challenges, we propose and implement C-VUE, a novel video encoder and the according training and inference system.
-C-VUE can perform real-time and prompt video understanding on streaming video in a continuous manner.
-That is to say, as each frame is received, the system updates its hidden state and outputs the events that have occurred so far, shown in Figure 1.
-Our C-VUE includes three key techniques.
-First, a long-range history modeling, designed to maintain the video historical context via the hidden state through a video-aware idea.
-Second, the temporal-relation guided spatial redundancy reduction, crafted to build a more efficient video modeling at the feature level.
-Third, a parallel training structure and a frame-weighted loss for the understanding of multi-scale events in the streaming video.
-Evaluation results.
-We evaluate C-VUE across various downstream tasks, comparing it with widely-used models for video understanding, including Transformer (vaswani2017attention, ), RNN (LSTM) (hochreiter1997long, ), CNN (kar2017adascan, ), as well as the SOTA video understanding system for edge scenarios, i.e., MoviNetA0 (kondratyuk2021movinets, ) and Mvitv2 (li2022mvitv2, ).
-The evaluation is conducted to demonstrate five capabilities: (i) the understanding of events with various temporal lengths; (ii) prompt recognition of an event as more frames are received; (iii) generality to long-tail events; (iv) the localization and understanding of continuous input events; (v) long-range memorization.
-The real-world video datasets used are FineGym (shao2020finegym, ) (sports gymnasium dataset), and Charades (sigurdsson2016hollywood, ) (daily indoor activities).
-We also compose more challenging datasets, e.g., concatenating all videos from the FineGym dataset into one to generate an hour-long video stream.
-Results show that with significantly fewer trainable parameters (ranging from 8.5k to 34.1M less than the baselines), C-VUE achieves accuracy improvements of 1.2% to 4.7% over Transformer, 2.6% to 5.9% over ConvPooling, 2% to 3% over LSTM, 14.5% to 67.4% over MoviNetA0, and 13.1% to 79.7% over MViTv2.
-Regarding latency, we evaluate C-VUE on typical edge devices (NVIDIA Jetson Orin, Apple MacBook, and an Intel PC with NVIDIA 4090 GPU).
-It achieves ¿30FPS on all these devices, up to 20$\times$ faster than the available inference systems.
-For the case study, we apply C-VUE video encoder to a video foundation model (Section 6).
-Specifically, we replace the visual encoder of the SOTA visual language model VideoLlama2 (damonlpsg2024videollama2, ) by C-VUE.
-The results demonstrate a 0.46-point improvement (on a 5-point scale) on the in-distribution dataset, and an improvement ranging from 1.19% to 4% on zero-shot datasets.
-Contributions. We summarize our contribution as follows:
-  • We propose a continuous video understanding system, C-VUE, based on adaptive state modeling. It promptly understands incoming video frames by maintaining historical context using a hidden state.
-  • C-VUE employs long-range history modeling to utilize the fixed-size hidden state for memorizing extensive historical data, reduces spatial redundancy based on temporal relations, and adopts a parallel training structure along with a frame-weighted loss function.
-  • Experiments show that C-VUE surpasses all baselines in continuous video understanding, using fewer parameters and achieving over 30 FPS on typical edge devices.
-A case study demonstrates that C-VUE can enhance the video foundation models on video QA tasks by serving as the video encoder.
-2. Background
-Continuous video understanding holds significant implications for complex human-computer interaction, embodied AI, and robots (duan2022survey, ; strafforello2023current, ; gadre2022continuous, ).
-While various approaches strive to achieve continuous understanding, none have solved this challenging task.
-All Frames at Once.
-(Figure 3(a)) This method performs the video understanding on all video frames, theoretically yielding the highest task accuracy (slowfast2019fei, ; kondratyuk2021movinets, ).
-However, its efficiency is a concern.
-A basic transformer model has complexity $\mathcal{O}(n^{2})$, making the whole process have $\mathcal{O}(n^{3})$ complexity.
-Applying optimized architecture can decrease it to $\mathcal{O}(n^{2})$.
+Take a Transformer model as an example, with the number of streaming frames $n$ grow, the computation cost of video activity recognition is $O(n^{3})$.
+Even with the many linear alternatives of Transformers, such as Linformer [41], RetNet [36], Mamba [13], and RWKV [28], the cost is still $O(n^{2})$.
 Our experiment (Figure 4) shows that as frames increase, FPS drops to 0.05 times, a 95% decrease.
 Fixed Window Size.
-(Figure 3(b)) This approach uses a fixed window size, like sliding window or sampling, to lower computational burden (Lin_2019_ICCV, ; Jiang_2019_ICCV, ; Li_2020_CVPR, ; li2022mvitv2, ).
-More efficient than the first method, its complexity ranges from $\mathcal{O}(mn)$ to $\mathcal{O}(m^{2}n)$, where $m$ equals the window size.
-The main challenge here is choosing an appropriate window size, as real-world events vary in duration.
-Small windows might miss longer events, and large ones could hinder understanding of shorter events.
+The most commonly used technique to control the cost is to fix the input size of the model even with the growing number of frames, such as sliding window or sampling [23, 17, 21, 22], shown in Figure 3(b).
+For example, TSN [40] first divides the past video into segments and selects a random snippet from each segment for activity recognition.
+VideoLlama2 [8] only samples 8 or 16 frames as input for any video length.
+This method has to ignore many detailed features.
+VideoMamba [20] adopt the Mamba [13] as the architecture to perform the efficient video understanding, however, it still adheres to the batch-based processing strategy.
+The main challenge is selecting an optimal window size to balance the detection of both short and long activities.
 State Assistance.
-(Figure 3(c)) By maintaining a group of hidden states, this method significantly enhances efficiency by reducing complexity to approximately $\mathcal{O}(n)$ (shi2015convolutional, ; huang2020efficient, ).
-However, as the input size grows, the fixed-size state struggles to remember the increasing history of inputs.
-Existing designs handle states in a video-unaware manner, preventing full utilization of the hidden states and the inability to model long-range history.
-Figure 4.
-Processing speed of understanding a 100k-frame video on RTX4090.
-Baselines 1(A) and 2(A) represent Transformer and Convpooling-based solutions using the “All Frames at Once” design.
-Baselines 1(F) and 2(F) employ the “Fixed window size” design (the window size is 100).
-Baseline 3 denotes an LSTM-based solution leveraging “State Assistance” design.
-3. C-VUE Design
-3.1. Design Overview
-Design Requirements.
-To continuously process the video stream, an ideal design should integrate the advantages of the aforementioned approaches.
-The first advantage is efficiency: input only the current frame instead of repeating or tracing back through previous frames (similar to the ”State Assistance” design).
-The second advantage is accuracy: have a comprehensive memory of historical information at all times (approximating the ”All Frames at Once” design).
-Design Challenges.
-To achieve the ideal design, the key lies in managing the growing volume of video, including: First, maximizing the memorization of the historical video data in a bounded subspace.
-Second, enabling processing at the edge in real-time without any increase in processing time as the video length extends.
-Design Preliminary.
-We base our method on a state-assistance design.
-The typical state-assistance design involves recurrent updates over a sequence through hidden states (sieber2024understanding, ).
-We take the RNN as an example to introduce the standard state-assistance design.
+There are also works [32, 15, 47], shown in Figure 3(c), that only input one current frame into a model (e.g., CNN) first to extract spatial features, and then conduct feature aggregation (e.g., through pooling [18] or LSTM [14]) among frames over time to capture temporal dependencies by hidden states.
+By maintaining a group of hidden states, this method significantly enhances efficiency by reducing complexity to approximately $\mathcal{O}(n)$ [32, 15].
+However, as the input size grows, the fixed-size state struggles to remember the increasing history of inputs [30, 40].
+Figure 5: Video frames of the balance beam sports, selected from the FineGym99 dataset [30]. The entire gymnastic movement was captured in only 17 frames, lasting less than 700 milliseconds.
+3 CARS Design
+A motivating video sample.
+Video streams contain both spatial (within-frame) and temporal (between-frame) information, both critical for accurate recognition.
+Figure 5 presents a video sample from a real-world gymnastics match to motivate our design and explain the challenges we faced.
+There are two issues we can observe from the video sample.
+Issue 1: the activity-related spatial feature is sparse.
+The activity feature in a video sample may be captured by only a small portion of the pixel content.
+Issue 2: object movement causes poor locality at the pixel level.
+It is hard to accurately recognize activities merely by analyzing two adjacent regions (by setting thresholds).
+We have observed the efficiency advantages of state-assistant methods in handling streaming video.
+Therefore, we utilize a state-assistant design, which involves recurrent updates over a sequence through hidden states [33].
+Our CARS has two key designs, i.e., activity spatial feature extraction and activity-aware state update, to solve these two issues.
+Figure 6 illustrates the system workflow.
+During the usage, streaming video is continuously input into the system frame by frame (without sampling).
+Our CARS extracts activity-related features and sends them to the activity-aware state update module.
+During the state update, the system attempts to identify possible activities and uses the activity features to update the corresponding part of the hidden state.
+If no further activity features related to those features maintained in the hidden state arrive, the corresponding video context is cleared from the hidden state.
+The updated state is stored in the system, waiting for the next frame to update it.
+Simultaneously, the updated hidden state can be sent to a classifier for timely activity recognition.
+With our optimization, the entire system can perform real-time video activity recognition on typical edge devices.
+Figure 6: CARS’s workflow for continuous activity recognition via adaptive video context modeling by selectively compressing only activity-related features into the hidden state.
+3.1 Activity Spatial Feature Extraction
+The challege of addressing issue 1 is that we lack supervised data to differentiate between redundant spatial features and those relevant to activities.
+Because the types of activities in the real world are diverse.
+We cannot merely retain an object’s pose and position.
+For instance, in Figure 5, it is crucial to also keep the feature of the balance beam present in the environment.
+In this way, we can accurately comprehend that this is a gymnastics activity, rather than a simple jump.
+The core idea of our feature extraction is to utilize the information bottleneck [38] as an unsupervised method to extract task-specific features.
+Specifically, we employ an information reduction module aimed at minimizing input features while maximizing task accuracy during training.
+Through this adversarial constraint, it can reduce redundancy and effectively achieve activity-related feature extraction.
+Formally speaking, let $X$, $X^{\prime}$, and $Z$ denote the random variables that represent all spatial features, extracted activity-related features, and redundant features, respectively.
+In our scenario, the activity spatial feature extraction determines how the feature $X^{\prime}$ is extracted from $X$ for the recognition task $f_{\theta}$.
+The goal of the activity-related feature extraction is to minimize the inclusion of redundancy information $Z$ in extracted feature $X^{\prime}$ while ensuring maximum data utility, i.e.,
+$\min_{\mathcal{I}(X^{\prime};Z)<\epsilon}\|f_{\theta}(X^{\prime}),f_{\theta}(X)\|,$ (1)
+[…]
+However, given the lack of supervised data regarding redundant information $Z$, it is hard to minimize the $\mathcal{I}(X^{\prime};Z)$.
+Note that, $\mathcal{I}(X^{\prime};Z)\leq\mathcal{H}(X^{\prime})$, where $\mathcal{H}$ is the entropy of a random variable.
+So, we have the relaxation of Equation 2, i.e.,
+$\min_{\mathcal{H}(X^{\prime})<\epsilon}\|f_{\theta}(X^{\prime}),f_{\theta}(X)\|.$ (2)
+That is, by minimizing the information of extracted features $X^{\prime}$ and preserving the accuracy of activity recognition, we can remove the redundant features as much as possible.
+Please note that the data utility objective, $\min\|f_{\theta}(X^{\prime}),f_{\theta}(X)\|$, is optimized by the accuracy of the activity recognition task (Equation 5).
+To reduce the information entropy of extracted features $X^{\prime}$, we propose two steps, i.e., the pooling layer and an information funnel Specifically, in our activity spatial feature extraction module, we first a pre-trained powerful feature extractor, i.e., CLIP [29], to extract general-purpose spatial features from an input frame.
+The CLIP model divides the input frame into multiple patches and computes an embedding for each patch.
+We utilize the average pooling to aggregate these embeddings by computing their mean, generating a single embedding.
+This constitutes the first stage of information reduction for $X^{\prime}$.
+After the first stage, each frame is reduced to a single embedding.
+Subsequently, this embedding is passed through our information funnel, which consists of a linear layer followed by an activation function, where the output size of the linear layer is smaller than its input size.
+This completes the second stage of information reduction for $X^{\prime}$.
+3.2 Activity-Aware State Update
+The challege of addressing issue 2 lies in online associating the features of each incoming frame with those in the video context, i.e., hidden state, thereby enhancing recognition accuracy.
+This state update serves as the association process since the recognition results are based on the updated state.
+We take the RNN as an example to introduce the traditionalstate update.
 It computes $\mathbf{y}$ through a dynamic recurrence of input signals $\mathbf{x}$ at each time step $t$ through hidden states $\mathbf{s}$.
 Formally,
-(1) $s_{t}=\phi_{h}(\mathbf{A}s_{t-1}+\mathbf{B}x_{t}+\mathbf{b}_{h}),$
-(2) $y_{t}=\phi_{o}(\mathbf{C}s_{t}+\mathbf{b}_{o}),$
+$s_{t}$ $=\phi_{h}(\mathbf{A}s_{t-1}+\mathbf{B}x_{t}+\mathbf{b}_{h}),$ (3)
+$y_{t}$ $=\phi_{o}(\mathbf{C}s_{t}+\mathbf{b}_{o}).$
 where $\mathbf{A},\mathbf{B}$, and $\mathbf{C}$ denote the state matrix, input matrix, and output matrix, respectively.
-The $\mathbf{b}_{h}$ and $\mathbf{b}_{o}$ are the bias parameters, and the functions $\phi_{h}$ and $\phi_{o}$ are activation functions.
-Subsequently, the gating mechanism is introduced to enhance the memory capability of state-assistance designs, e.g., LSTM (sherstinsky2020fundamentals, ).
-More specifically, gates denoted as $g^{s}_{t}$, $g^{i}_{t}$, and $g^{o}_{t}$ are introduced at each state, input, and output, respectively.
-At time $t$, these gates dynamically adjust the state update process and the output based on the given input.
+The $\mathbf{b}$ denotes the bias parameters, and the functions $\phi$ denotes activation functions.
+However, the RNN and its gating-based variants, e.g., LSTM [31], are widely considered challenging to maintain input context by many research works [12, 45, 27].
+Our experimental results indicate that the LSTM-based baseline performs poorly on streaming video inputs (Table 2).
+We argue that the underperformance of such methods is due to their inefficient utilization of the hidden state.
+Specifically, standard RNN processes input at different moments identically, i.e., fixed $A$, $B$, and $C$ in Equations 3, leading to irrational use of the hidden state (Figure 2).
+We argue that this deficiency stems from the state update method’s inability to adapt effectively to incoming video frames, as it applies one update strategy to all frames, irrespective of their distinct features.
+We enhance the state update method by enabling activity-aware hidden state updates based on activity spatial features.
+The core idea is to introduce dynamic adaptability, akin to the attention mechanism in transformers, which adjusts its focus across input sequences to accommodate varying inputs.
+Specifically, we replace the fixed matrix used for state updates with an input-dependent update manner.
+Our input-dependent update takes the current activity spatial features as input and determines how the current input should be incorporated into the hidden state.
 Formally,
-(3) $s_{t}=\phi_{h}(g^{s}_{t}\circ\mathbf{A}s_{t-1}+g^{i}_{t}\circ\mathbf{B}x_{t}+\mathbf{b}_{h}),$
-(4) $y_{t}=\phi_{o}(g^{o}_{t}\circ\mathbf{C}s_{t}+\mathbf{b}_{o}).$
-This gating mechanism resembles a weighting scheme, represented as a vector with values ranging from 0 to 1, which is computed based on the input and the last output.
-Limitation of RNN and Its Variants.
-The RNN and its gating-based variants are widely considered difficult to train and challenging to handle long-range historical data by many research works (greaves2019statistical, ; zhao2020rnn, ; miller2018stable, ).
-Our experimental results also indicate that the LSTM-based baseline performs poorly on hours-long videos (Table 2).
-We argue that the underperformance of such methods is due to their inefficient utilization of the hidden state.
-Specifically, standard RNN processes input at different moments identically, i.e., fixed $A$, $B$, and $C$ in Equations 1 and 2, resulting in a lack of adaptation towards these inputs.
-When redundancies exist within the input data, this leads to an irrational use of the hidden state.
-Despite the subsequent introduction of gating mechanisms to enhance input adaptation, these designs merely adopt a weighted approach to decide how the new inputs should merge with the existing states.
-Such a design idea lacks flexibility and fails to fully tap into the memorization capability of the hidden state.
-Our Key Designs.
-We introduce our C-VUE with the following three parts.
-First, long-range history modeling (Section 3.2).
-This is our core design, aiming to maximize the memorization capability of multi-scale events over long sequences.
-This design adopts two techniques: 1) a fully video-aware mechanism to improve the hidden state’s utilization; and 2) a scale invariance mechanism for multi-scale events understanding.
-Second, temporal relation-guided spatial redundancy reduction (Section 3.3).
-Performing history modeling in pixel space can result in excessive system overhead and increased learning difficulty.
-This design enables the efficient modeling by eliminating the redundant information at the feature level with the guidance of the temporal relation.
-Third, parallal training with frame-weighted loss (Section 3.4).
-It introduces how our C-VUE builds the training structure for continuous long video understanding.
-We detail the efficient parallel training for long video samples and the weighted loss design for the understanding of every frame.
-3.2. Long-range History Modeling
-The challenge of understanding long-range history primarily stems from two aspects, namely inter-event noise and intra-event sparsity.
-The inter-event noise implies that there may be numerous noise frames between two events.
-In the context of continuous understanding, the longer the video, the more pronounced this issue becomes.
-The intra-event sparsity represents the inherent feature sparsity of event information.
-Some events carry limited information despite their extended durations, e.g., sleeping.
-However, the multi-scale characteristics of events make it difficult to handle this sparsity.
-To handle these issues, we propose two techniques.
-The first is a fully video-aware history modeling.
-Here, each fixed parameter is improved to a dynamically-generated parameter based on the input, enhancing the ability to filter noise between events.
-The second is a scale-invariant event memorization technique, which dynamically determines how an event should be memorized in the hidden state based on its information density, rather than its duration.
-3.2.1. Fully Video-aware History Modeling
-In existing designs, once trained, the model parameters cannot be flexibly adjusted based on the input frame.
-Intuitively, if some frame is repetitive or does not contain any events, the design should discard this frame without updating the hidden state, and simply copy the result from the last output.
-Therefore, we aim to improve all matrix parameters into input-dependent ones.
-Our design incorporates dynamic adaptability, similar to the attention mechanism in transformers.
-The attention mechanism adapts to different inputs by adjusting its focus on various parts of the input sequence.
-Many efforts have explored that, to model the long-range history, the matrix $A$ should be designed carefully.
-So, in our design, we only improve matrics $B$ and $C$ to input-dependent ones.
-Then we detail how to improve the matrix $A$ for long-range history modeling.
-We improve the $B$ and $C$ as functions of the input, i.e., $B_{t}=f_{B}(x_{t})$ and $C_{t}=f_{C}(x_{t})$.
-Consequently, during training, we no longer learn the $B$ and $C$ matrices directly but instead learn how to dynamically generate these matrices based on the input.
-The hidden state is maintained dependent on the input $x$ to only keep the event features.
-In our method, we utilize the two linear layers as the function $f_{B}$ and the function $f_{C}$, respectively.
-Matrix $A$ controls how information from previous states is transferred to the new state.
-The design of matrix $A$ determines whether the method retains only recent states (local dependency) or captures all past states.
-Recently, many efforts have been put into how to improve the ability to memorize long-range history by designing matrix $A$.
-Inspired by Work (gu2023mamba, ), we redesign the transform matrix $A$ and its initialization method.
-First, we enhance matrix $A$ by making it a diagonal matrix, which has been proven to significantly improve the capability to model long-range history.
-Various efforts have explored alternative initialization techniques for matrix $A$ during training to boost model performance.
-Following this work, we employ random initialization techniques, which achieve similar results.
-Therefore, in our C-VUE, we utilize a randomly initialized diagonal matrix $A$ to enable long-range memorization.
-3.2.2. Scale-invariance Events Memorization
-As shown in Figure 2, different events contain varying numbers of frames, but they may convey a similar amount of information.
-We aim to use hidden states of similar sizes to memorize events with the same amount of information, even though these events have different time scales.
-Inspired by Work (nguyen2022s4nd, ), we can leverage discretization techniques to dynamically resample video frames.
-The input frames $x_{0}$, $x_{1}$, …, $x_{t}$ can be viewed as uniformly sampled from an underlying signal $f(x)$ with a step size $\Delta$, i.e., $x_{t}=f(t\Delta)$.
-The standard RNN utilizes matrics $A$ and $B$ to update the state in keeping with $\Delta$ (Equation 1 and 2).
-To achieve our design goal, we no longer use video frames as the basic unit for sampling but instead use the events as the basic unit for sampling via a dynamic $\Delta$.
-Then, we apply the dynamic $\Delta$ on the state update procedure to achieve the dynamic resampling.
-Note that the step size $\Delta$ is not a real sampling rate, and we do not perform the resampling technique on the video frames.
-All input frames are processed during the continuous video understanding procedure.
-We utilize the discretization technique, i.e., zero-order hold (ZOH), utilized in (gu2023mamba, ) to achieve the above dynamic resampling.
-Specifically, $A$ and $B_{t}$ can be improved with an input-dependent parameter $\Delta_{t}=f_{\Delta}(x_{t})$:
-(5) $\overline{A_{t}}=\exp(\Delta_{t}A)$
-(6) $\overline{B_{t}}=(\Delta_{t}A)^{-1}\exp(\Delta_{t}A-I)\cdot\Delta_{t}B_{t}$
-$\Delta_{t}$ is calculated based on the input $x_{t}$, thus achieves the awareness of the current event. We use a linear layer as the $f_{\Delta}$ function.
-3.3. Temporal Relation-guided Spatial Redundancy Reduction
-Videos contain both spatial (within-frame) and temporal (between-frame) information, both critical for accurate understanding.
-However, as shown in Figure 5, video data contains a vast amount of redundant information in spatial space.
-Directly performing history modeling introduces lots of redundant computations, leading to poor system performance.
-Eliminating the redundancy is challenging.
-Because, firstly, the event feature is sparse.
-The event feature in a video sample may be captured by only a small portion of the pixel content.
-Secondly, object movement causes poor locality at the pixel level.
-It is hard to accurately understand events merely by analyzing two adjacent regions (by setting thresholds).
-Therefore, if we perform video understanding at the pixel level like traditional approaches, e.g., MoviNet (kondratyuk2021movinets, ) and Mvitv2 (li2022mvitv2, ), there are lots of redundant computation (as shown in the left part of Figure 6).
-Figure 5.
-Video frames of the balance beam sports, selected from the FineGym99 dataset (shao2020finegym, ).
-The entire gymnastic movement was captured in only 17 frames, lasting less than 700 milliseconds.
-Figure 6.
-Different ideas to model the video information.
-Traditional methods associate any two pixels or patches with complexity $\mathcal{O}(n^{2})$ or $\mathcal{O}(m^{2})$, where $n$ and $m$ are the pixel and patch counts.
-By reducing spatial redundancy, we model video content more efficiently.
-Given the issues mentioned earlier, many pixels, such as background pixels in Figure 5, do not contain event information, leading to computational redundancy.
-Therefore, such a method inevitably leads to lots of computation redundancy.
-Additionally, due to low inter-frame pixel locality caused by object movement, it is challenging to reduce the quadratic complexity simply by searching nearby pixels.
-Traditional video processing methods have carried over into the era of LLM-based video foundation models (damonlpsg2024videollama2, ).
-The vision encoder break a frame into patches and then correlate these patches across frames in an all-to-all manner.
-The patch number is $m$, and the complexity is $\mathcal{O}(m^{2})$.
-This approach makes it difficult for video foundation models to balance accuracy and system performance.
-Correlating a large number of frames is costly in both the training and inference phases.
-Specifically, the high computational complexity often requires excessive downsampling, which reduces accuracy.
-We analyze this issue in Section 6 and show how our design can improve video foundation models.
-Our approach introduces a temporal relation-guided spatial redundancy reduction.
-The key insight is that we can learn the rule to reduce the redundant spatial information, that is not related to the event, by correlating these features in the temporal dimension.
-For example, understanding a jumping action, as shown in Figure 5, is primarily about observing how the person’s pose and position change between frames, which is unrelated to what the person is wearing.
-Hence, we can first attempt to capture the specific pose and position information of the objects at each moment, then associate them over time for further understanding.
-However, it is challenging, as the types of events in the real world are diverse.
-We cannot merely retain an object’s pose and position.
-For instance, in Figure 6, it is crucial to also keep the feature of balance beam present in the environment.
-In this way, we can accurately comprehend that this is a gymnastics action, rather than a simple jump.
-To achieve the reduction goal, we first use a pre-trained powerful feature extractor to extract general-purpose spatial feature, and then use a lightweight projection layer to perform the reduction.
-Specifically, we utilize the CLIP (radford2021learning, ) to perform the preliminary spatial feature extraction.
-CLIP is one of the most powerful vision encoders, which serves as the vision encoder in mainstream video foundation models (awais2023foundational, ; madan2024foundation, ).
-Then we utilize a linear layer followed by an activation function as the projection layer to reduce the redundancy all-purpose spatial features.
-The linear layer is trained together with our history modeling layer to reduce the redundancy with the guidance of the temporal relationship.
-After the reduction, the spatial feature can be used to enable efficient long-range history modeling, as shown on the far right of Figure 6.
-Thus, the complexity of modeling the relationship between two frames is reduced to $\mathcal{O}(1)$.
-3.4. Parallel Training with Frame-weighted Loss
-Given the recurrent form of our C-VUE may significantly affect the system performance during the training phase.
-We adopt a parallel training structure to efficiently train long video samples.
-Additionally, we design a frame-weighted loss to calculate the loss for each frame, thereby accomplishing prompt video understanding.
-3.4.1. Efficient Parallel Training
-To implement the redundancy reduction design described in Section 3.3, we start by using a pre-trained vision encoder ($\theta_{\texttt{spatio\_encoding}}$) to extract the general-purpose spatial feature.
-It is important to note that this vision encoder remains frozen during training.
-Before performing history modeling, we project the frame embeddings using a linear layer with ReLU activation ($\theta_{\texttt{reduction}}$).
-Recall that this projection is used to perform spatial redundancy reduction based on the temporal relationship.
-The $\theta_{\texttt{history\_modeling}}$ is the design proposed in Section 3.2 for adaptive history modeling.
-We unfold the recurrent state transition process into a convolutional form for parallel computation.
-Finally, we employ another linear layer with ReLU activation ($\theta_{\texttt{post\_proj}}$) to map the output of each time step to the space of the downstream task.
-The overall structure of our model is illustrated on the right side of Figure 7.
-The training process is detailed in Algorithm 1.
-Although our system (C-VUE) employs a recurrent design, processing one frame at a time.
-We achieve high parallelism during training.
-Specifically, in the vision encoder, as well as during spatial feature reduction and post processing, we treat the temporal dimension as a batch dimension for parallel computation.
+$A\rightarrow A_{t}=f_{A}(x_{t});B\rightarrow B_{t}=f_{B}(x_{t});C\rightarrow C_{t}=f_{C}(x_{t}).$ (4)
+That is, during training, we no longer learn the $A$, $B$, and $C$ matrices directly but instead learn how to dynamically generate these matrices based on the input.
+Here, we employ Mamba [13] as the implementation of our activity-aware state update module, as it aligns with our design philosophy: it dynamically computes these matrices based on each input.
+Additionally, Mamba’s implementation accelerates the training process for such recursive formulations.
+Note that while Mamba has been explored for video processing [20, 8, 7], these works adopt strategies such as processing “all frames at once” or using a “fixed window size”, treating Mamba primarily as a lightweight alternative to Transformers.
+In contrast, our “state-assistant” design fundamentally differs in both approach and purpose.
+3.3 Training Structure
+The training structure is detailed in Algorithm 1.
+The CARS’s architecture consists of a pre-trained vision encoder ($\theta_{\texttt{CLIP}}$), a feature extractor ($\theta_{\texttt{extractor}}$), an activity-aware state update module ($\theta_{\texttt{state\_maintaince}}$), and a classifier ($\theta_{\texttt{cls}}$).
+The vision encoder extracts the general-purpose spatial feature and is frozen during training.
+The feature extractor is a linear layer with ReLU activation.
+The activity-aware state update module is a single Mamba layer.
+The classifier is a ReLU activation.
 Data: Video frames and corresponding labels: $\texttt{X}_{1:n}$ and $\texttt{Y}_{1:n}$
@@
 3      while not convergent do
-             $\texttt{Emb}_{1:n}\leftarrow\theta_{\texttt{spatio\_encoding}}^{\star}(\texttt{X}_{1:n})$ // parallel
-             $\texttt{Tok}_{1:n}\leftarrow\theta_{\texttt{reduction}}(\texttt{Emb}_{1:n})$ // parallel
-             $\texttt{Ir}_{1:n}\leftarrow\theta_{\texttt{history\_modeling}}(\texttt{Tok}_{1:n})$ // in convolution mode
-             $\hat{\texttt{Y}}_{1:n}\leftarrow\theta_{\texttt{post\_proj}}(\texttt{Ir}_{1:n})$ // parallel
+4             $\texttt{Emb}_{1:n}\leftarrow\theta_{\texttt{CLIP}}^{\star}(\texttt{X}_{1:n})$
+5             $\texttt{Tok}_{1:n}\leftarrow\theta_{\texttt{extractor}}(\texttt{Emb}_{1:n})$
+6             $\texttt{Ir}_{1:n}\leftarrow\theta_{\texttt{state\_maintaince}}(\texttt{Tok}_{1:n})$
+7             $\hat{\texttt{Y}}_{1:n}\leftarrow\theta_{\texttt{cls}}(\texttt{Ir}_{1:n})$
 8             $\texttt{L}\leftarrow\mathcal{L}(\texttt{Y}_{1:n},\hat{\texttt{Y}}_{1:n})$
 9             $\theta\leftarrow\theta-\nabla_{\theta}\texttt{L}$
-6
+10
 11      return $\theta$
-Algorithm 1 The parallel training structure of C-VUE. $\theta_{\texttt{spatio\_encoding}}^{\star}$ is frozen during the training.
-Figure 7. The model architecture of our C-VUE.
-3.4.2. Frame-weighted Loss for Continuous Understanding
-Here we introduce the loss function $\mathcal{L}$ in Algorithm 1.
-Video datasets are often categorized as either motion-focused or scene-focused ones.
-The former requires understanding the process of action transitions to interpret the video accurately.
-For instance, in gymnastics sports datasets (shao2020finegym, ), the initial frames of turning once and twice may look similar, but the later frames are crucial for determining the action category.
-In contrast, scene-focused tasks can often be distinguished by certain static appearances, such as drinking water or reading (sigurdsson2016hollywood, ).
-Video datasets are usually annotated in a temporal-grounded manner.
-To account for the different characteristics of the video dataset, we propose a frame-weighted loss.
-The key principle is to compute the loss by taking a weighted average across all frames.
-Formally, this can be expressed as:
-(7) $\mathcal{L}_{v}=W^{T}\times[\mathcal{L}_{f}^{1},\mathcal{L}_{f}^{2},\ldots,\mathcal{L}_{f}^{N}],$
+Algorithm 1 The parallel training structure of CARS. $\theta_{\texttt{spatio\_encoding}}^{\star}$ is frozen during the training.
+The loss function $\mathcal{L}$ is a frame-weighted loss.
+The core intuition is that activity recognition can be categorized into two types: motion-focused (requiring temporal information, e.g., performing a triple somersault) or scene-focused (relying solely on static information, e.g., drinking water).
+Formally, it can be expressed as:
+$\mathcal{L}_{v}=W^{T}\times[\mathcal{L}_{f}^{1},\mathcal{L}_{f}^{2},\ldots,\mathcal{L}_{f}^{N}],$ (5)
 where $W$ is the weight vector of the loss from Frame 1 to Frame $N$.
-Users can design different weight vectors based on their datasets.
-For motion-focused datasets, we recommend a linear growth strategy, where the weight of the loss for the initial frame is relatively low and gradually increases over time within the video frames.
+For motion-focused tasks, we recommend a linear growth strategy, where the weight of the loss for the initial frame is relatively low and gradually increases over time within the video frames.
 Specifically, $W=[\frac{i}{N^{2}}]_{i=1}^{N}$.
@@
 That is, $W=[\frac{1}{N}]_{i=1}^{N}$.
-4. Implementation
-Network Architecture.
-We implement the $\theta_{\texttt{spatio\_encoding}}^{\star}$ with the powerful vision encoder, i.e., CLIP (clip_impl, ).
-$\theta_{\texttt{temporal\_modeling}}$ is implemented with a single layer Mamba (gu2023mamba, ).
-Here we utilize Mamba implementation for acceleration.
-Mamba proposes a selective scanning algorithm, which assumes the order in which it does operations does not matter through the associate property.
-It calculates the sequences in parts and iteratively combines them to calculate the history modeling in a convolution mode.
+4 Evaluation
 Hyperparameter Setting.
-During the training, we freeze the vision encoder and only train the two projection layers and the Mamba layer.
-The AdamW (loshchilov2017decoupled, ) is adopted as the optimizer.
+The AdamW [24] is adopted as the optimizer.
 The learning rate changes dynamically, using Cosine annealing with warm restarts.
@@
 Optimization Deployment.
-To ensure real-time operation and cross-platform deployment, we have utilized compilation and quantization techniques on our model.
-We have encoded the model file into $C$ code to guarantee cross-platform deployment, and ultimately achieved deployment on a MacBook Air (M3, 2024), Jetson AGX Orin 64 GB, and a PC with an RTX 4090 GPU card.
-We implemented our vision encoder CLIP (radford2021learning, ) based on the llama.cpp framework (llama_cpp_impl, ), completed the compilation, and carried out INT8 quantization.
-The existing CLIP.cpp (clip_cpp_impl, ) framework based on llama.cpp for compiling CLIP only supports CPU backend, hence it can’t utilize the GPU on mobile devices.
-Other mainstream frameworks lack good compatibility for cross-platform deployment.
-For example, OpenClip (open_clip_impl, ) and ONNXRuntime-GPU (onnx_impl, ) lack support for using the CLIP model with Jetson GPU.
-We have expanded support for ggml in our implementation, enabling GPU-based inference on all the aforementioned platforms, which resulted in significant system performance improvement.
-Please refer to Figure 9.
-The official implementation of Mamba (mamba_impl, ) only supports the PyTorch framework with the CUDA backend.
-We have compiled our Mamba part into a $C$ program, thereby allowing its deployment on various platforms.
-5. Evaluation
-5.1. Experiment Setups
-Evaluation Goals. We aim to evaluate five capabilities of the system:
-Capability 1: Understanding of events with various lengths.
-Capability 2: Prompt response as more frames are received.
-Capability 3: Generality for long-tail data.
-Capability 4: Understanding of continuous input events.
-Capability 5: Long-range memorization.
-Figure 8.
-Datasets used in experiments.
-S represents a sample; a line segment represents the duration of an event; E represents a label.
-In Charades and Multi-label FineGym, a frame may have multiple labels.
+We implemented the CLIP vision encoder using the llama.cpp framework [2], performing INT8 quantization for improved efficiency.
+While the CLIP.cpp framework [1], based on llama.cpp, only supports CPU backends, limiting GPU use on mobile devices, other frameworks like OpenClip [5] and ONNXRuntime-GPU [4] also lack support for Jetson GPUs.
+We extended support for ggml format, enabling GPU-based inference across all platforms, which resulted in a significant performance boost (see Figure 9).
+Furthermore, we compiled the Mamba implementation [3], which originally supported only PyTorch with CUDA, into ggml format for cross-platform deployment.
+Through these optimizations, we successfully deployed our CARS on a MacBook Air (M3, 2024), Jetson AGX Orin 64GB, and a PC with an RTX 4090 GPU.
+Figure 7: Datasets used in experiments. S represents a sample; a line segment represents the duration of an activity; E represents a label. In Charades, a frame may have multiple labels.
+4.1 Experiment Setups
+Evaluation Goals. We aim to evaluate four capabilities:
+Capability 1: Recognition of activities across timescales.
+Capability 2: Activity Recognition in streaming input.
+Capability 3: Prompt response as more frames are received.
+Capability 4: Generality for long-tail activities.
 Datasets.
 We utilize five datasets to evaluate different capabilities.
-FineGym99 (shao2020finegym, ) (For capabilities 1 & 2).
+FineGym99 [30] (For capabilities 1 & 3).
 It is a gymnasium video dataset.
 Each video clip is labeled with a gymnastics movement.
-The duration of events varies from 2 to 50 seconds.
-FineGym288 (shao2020finegym, ) (For capabilities 1, 2, & 3). It is an extended version of FineGym99, which enhances its by adding categories. FineGym288 has the natural long-tail distribution.
-Charades (sigurdsson2016hollywood, ) (For capabilities 1, 2, & 5).
+The duration of activities varies from 2 to 50 seconds.
+FineGym288 [30] (For capabilities 1, 3, & 4).
+It is an extended version of FineGym99, which enhances its by adding categories.
+FineGym288 has a natural long-tail distribution.
+Charades [34] (For capabilities 1 & 3).
 It is a multi-label dataset, composed of daily indoor activities with an average length of 30 seconds.
 Each video clip consists of about 800 frames on average, varying from 32 to 3,944 frames.
-We have re-annotated the video samples for continuous video understanding as Figure 8(b).
-All-event-in-one FineGym (For capabilities 1, 2, & 4).
+We re-annotated Charades, as Figure 7(b), to evaluate the method’s effectiveness in maintaining video context when multiple activities occur simultaneously.
+All-activity-in-one FineGym (For capabilities 1, 2, & 3).
 It is constructed based on the FineGym99 as shown in Figure 7(c).
 We concatenated all the video clips from FineGym99’s testing dataset, forming a single long video lasting 3.5 hours and consisting of over 440k frames.
-Multi-Label FineGym@k (For capabilities 1, 2, & 5).
-It is constructed based on the FineGym99 as shown in Figure 8(d).
-The new video clips are constructed by randomly concatenating $k$ events of FineGym99 together and re-annotating them.
+The concatenated video is employed to simulate streaming video containing continuous activities.
 Baseline Methods.
 We utilize five baselines to execute continuous perception, whereby the method predicts an outcome at each timestep.
-Baselines 1, 2, and 3 are constructed following different video processing paradigms.
-Baselines 4 and 5 are state-of-the-art (SOTA) methods for balancing accuracy and efficiency for video classification.
-Baseline 1: Transformer (vaswani2017attention, ).
-The overall architecture mirrors ours for a fair comparison, except that we replace our history modeling layer with a transformer of equivalent parameter size.
-Baseline 2: ConvPooling (kar2017adascan, ).
-The overall architecture mirrors ours for a fair comparison, except that we replace the history modeling layer with a convPooling of equivalent parameter size.
-The convPooling consists of a 1D convolution and a 1D adaptive max pooling.
-Baseline 3: LSTM (hochreiter1997long, ).
-This design keeps the overall structure consistent with ours for a fair comparison, substituting the history modeling layer with an LSTM of equal parameter magnitude.
-Baseline 4: MoviNetA0 (kondratyuk2021movinets, ).
+Baselines 1, 2, and 3 are constructed following different video processing paradigms (as shown in Figure 8).
+Their overall architecture mirrors ours, except that we replace the mamba layer with other networks.
+Baseline 1 utilizes a Transformer [39].
+Baseline 2 utilizes a convPooling [18], which consists of a 1D convolution and a 1D adaptive max pooling.
+Baseline 3 utilizes an LSTM [14].
+Baselines 4 and 5 are state-of-the-art (SOTA) methods for balancing accuracy and efficiency for video classification for on-device computing scenarios.
+Baseline 4: MoviNetA0 [19].
 This is a Mobile Video Network specifically designed for efficient video recognition.
@@
 Drawing from the experience of using this model on the Charades dataset, we similarly down-sample the video to 6 fps for video preprocessing.
-Baseline 5: Mvitv2(li2022mvitv2, ) This is an improved multiscale vision transformer for both classification and detection tasks.
+Baseline 5: Mvitv2[22] This is an improved multiscale vision transformer for both classification and detection tasks.
 This method necessitates the sampling of any given video to a fixed number of frames for video classification tasks.
 We trained the baseline that takes as input 16 frames at a time, which is one of the default settings.
+Figure 8: The constructed baselines 1, 2, and 3.
 Input Strategy. We utilize three typical input strategies to perform the evaluation.
-S1: All frames feeding. All frames from frame 1 up to frame $N$ are fed into the baseline method to provide the necessary context for achieving continuous perception at timestep $N$.
-S2: Sliding windowed feeding. All frames between the $(N-m)$-th frame to the $N$-th frame are fed into the baseline at timestep $N$, where $m$ denotes the sliding window size.
-S3: Single frame feeding. Only frame $N$ is supplied to the baseline method.
+S1: All frames feeding.
+All frames from frame 1 up to frame $N$ are fed into the baseline method to provide the necessary context for achieving continuous perception at timestep $N$.
+S2: Sliding windowed feeding.
+All frames between the $(N-m)$-th frame to the $N$-th frame are fed into the baseline at timestep $N$, where $m$ denotes the sliding window size.
+S3: Single frame feeding.
+Only frame $N$ is supplied to the baseline method.
 Metrics.
 In this work, we utilize various metrics, namely Accuracy, Mean Average Precision (mAP), and the Jaccard Index for single-label and multi-label video classification tasks.
-Furthermore, we introduce a novel metric referred to as the Early Detection Rate (EDR), designed to evaluate the model’s ability to detect events promptly within videos.
+Furthermore, we introduce a novel metric referred to as the Early Detection Rate (EDR), designed to evaluate the model’s ability to detect activities promptly within videos.
 The definition is $\mathtt{EDR}={(E_{pred}-E_{gt})}/{E_{N}}$, where $E_{pred}$ refers to the frame index at which the label is detected.
-$E_{gt}$ represents the ground truth frame index, that is, the index of the frame where the event begins.
-$E_{N}$ signifies the total number of frames in the event.
-Table 1.
-The accuracy and EDR results of different baselines.
+$E_{gt}$ represents the ground truth frame index, that is, the index of the frame where the activity begins.
+$E_{N}$ signifies the total number of frames in the activity.
+Table 1: The accuracy and EDR results of different baselines.
 The #Param counts the parameters that need to be trained.
@@
 Methods #Param Acc (%) EDR (%) Acc (%) EDR (%) $\text{mAP}_{\text{all}}$ (%) $\text{mAP}_{\text{last}}$ (%) EDR (%)
-Baseline 1 (Transformer) 210.5K 82.3 48.20 76.6 53.20 17.2 21.4 52
-Baseline 2 (ConvPooling) 367.3K 82.6 38.40 75.4 41.40 15.8 21 43.3
-Baseline 3 (LSTM) 218K 84.1 23.40 78.4 25.80 15.33 17.8 84
-Baseline 4 (MoviNet) 2.1M 19.7 41.60 13.9 44.20 3.9 5.8 32
-Baseline 5 (Mvitv2) 34.3M 6.4 20.50 3.1 24.10 5.3 7.3 99.9
+B1 (Transformer-based) 210.5K 82.3 48.20 76.6 53.20 17.2 21.4 52
+B2 (ConvPooling-based) 367.3K 82.6 38.40 75.4 41.40 15.8 21 43.3
+B3 (LSTM-based) 218K 84.1 23.40 78.4 25.80 15.33 17.8 84
+B4 (MoviNet) 2.1M 19.7 41.60 13.9 44.20 3.9 5.8 32
+B5 (Mvitv2) 34.3M 6.4 20.50 3.1 24.10 5.3 7.3 99.9
 Ours 202K 86.1 22.50 81.3 25.10 18.4 22.1 56.3
-5.2. Evaluation of All Five Capabilities
+4.2 Evaluation of All Capabilities
 In Table 1, we present the outcomes from several baseline models across different datasets.
@@
 For batch-based methods, i.e., Baselines 1, 2, 4, and 5, we adopted the all-frame feeding strategy.
-Although this significantly increases computational overhead, it enables these methods to have sufficient context to understand the video content.
+Although this significantly increases computational overhead, it enables these methods to have sufficient context to recognize the video activities.
 Both Baseline 3 and our method utilize a streaming processing paradigm, for which we selected the single-frame feeding strategy.
-In the evaluation of single-label classification datasets, we utilize Accuracy (Acc) as the primary measure of correctness.
+In evaluating single-label classification datasets, we utilize Accuracy (Acc) as the primary measure of correctness.
 This accuracy is ascertained by verifying whether the correct action is detected within a given video.
@@
 If a method fails to detect a particular label, this failure is taken into account during computation by adding a 100% increase to the result for that category.
-This signifies that the correct label was not detected even upon the conclusion of the video.
+This signifies that the correct label was undetected even after the video’s conclusion.
 Below we present our experimental analyses and the subsequent conclusions:
 Superior Accuracy: Our method consistently outperforms all baseline methods in terms of accuracy across all tasks, improving metrics by ranges of 2 to 79.7, 2.9 to 78.2, 1.2 to 14.5, and 0.7 to 16.3.
-The leading accuracy demonstrates that our C-VUE can effectively understand events of different scales (Capability 1).
-The superior accuracy on the FineGym288 dataset underscores our method’s generalization ability for long-tailed data (Capability 3)).
-Prompt Event Detection: Our approach exhibits an outstanding ability to promptly detect events (Capability 2) and continuously processing capability (Capability 4) by EDR.
-For single-label tasks, events are detected within the first quarter of their occurrence.
-For complex multi-label tasks, all events are identified halfway through the video.
-Memorization Capability: Within the multi-label dataset Charades, the $\text{mAP}_{\text{last}}$ denotes the ability to recognize all events in the video by its conclusion.
+The leading accuracy demonstrates that our CARS can effectively recognize activities of different scales (Capability 1).
+The superior accuracy on the FineGym288 dataset underscores our method’s generalization ability for long-tailed data (Capability 4)).
+Prompt Recognition : Our approach exhibits an outstanding ability to promptly detect activities (Capability 3) and continuously process (Capability 2) by EDR.
+For single-label tasks, activities are detected within the first quarter of their occurrence.
+For complex multi-label tasks, all activities are identified halfway through the video.
+Optimal Use of Trainable Parameters: Notably, our method retains superior performance while utilizing a minimal number of training parameters, amounting to only 202K.
+This represents a drastic reduction when compared to the parameter requirements of baseline models, with differences ranging from 8.5K to 34.1M.
+Robust Stability Across Varied Tasks: Our method showcases robust stability in tackling diverse tasks.
+It performs uniformly well across differing task scales (including those with varied activity durations), varied video characteristics (motion-focused and scene-focused video), and differing classification objectives (such as single-label and multi-label classifications).
+No baseline methods are currently able to match this level of comprehensive performance.
+Maintaining Video Context for Multiple Activities: Within the multi-label dataset Charades, the $\text{mAP}_{\text{last}}$ denotes the ability to recognize all activities in the video by its conclusion.
 Notably, Baselines 1 and 2 attained such accuracy by taking as input all video frames simultaneously.
@@
 Notably, our approach achieves 1.24 times the accuracy of Baseline 3, which employs LSTM for continuous processing.
-Results show that our C-VUE shows a better long-range memorization ability (Capability 5).
-Optimal Use of Trainable Parameters: Notably, our method retains superior performance while utilizing a minimal number of training parameters, amounting to only 202K.
-This represents a drastic reduction when compared to the parameter requirements of baseline models, with differences ranging from 8.5K to 34.1M.
-Robust Stability Across Varied Tasks: Our method showcases robust stability in tackling diverse tasks.
-It performs uniformly well across differing task scales (including those with varied event durations), varied video characteristics (motion-focused and scene-focused video), and differing classification objectives (such as single-label and multi-label classifications).
-No baseline methods are currently able to match this level of comprehensive performance.
+Results show that our CARS shows a better utilization of the hidden state.
 Considering the inadequate accuracy demonstrated by Baselines 4 and 5, subsequent experimental evaluations will only involve Baselines 1, 2, and 3, along with ours.
-Table 2. Evaluation results on All-event-in-one FineGym. (The parenthesized number of S2 signifies the window size, while 725 refers to the maximum frame count of a single event.)
+Table 2: Evaluation results on All-activity-in-one FineGym. (The parenthesized number of S2 signifies the window size, while 725 refers to the maximum frame count of a single activity.)
 Method Baseline 1 Baseline 3
@@
 Acc (%) 4.7 11.8 13.4 11.3 7 14.8
-5.3. Further Evaluation of Capabilities 1 & 4
-We assess the continuous perception capability with the All-event-in-one FineGym dataset, a 3.5-hour video described in Section 5.1.
+Figure 9: Processing speed of streaming video activity recognition with 10k frames on different platforms.
+4.3 Further Evaluation of Capabilities 1 & 2
+We assess the continuous perception capability with the All-activity-in-one FineGym dataset, a 3.5-hour video described in Section 4.1.
 All baseline methods use the model trained on the FineGym99 dataset, whose accuracy evaluated on a single video clip is reported in Table 2.
@@
 For batch processing approaches (Baselines 1 and 2), we test all three strategies.
-Regarding streaming processing methods (our C-VUE and Baseline 3), we use the single frame feeding strategy (Strategy 3).
-Both our C-VUE and Baseline 3 benefit from the model’s memorization of past video frames.
+Regarding streaming processing methods (our CARS and Baseline 3), we use the single frame feeding strategy (Strategy 3).
+Both our CARS and Baseline 3 benefit from the model’s memorization of past video frames.
 During continuous perception, each frame is processed only once without redundant input.
-The Acc metric is calculated by considering predictions of all frames in the video.
+The Acc metric is calculated by considering predictions of all video frames.
 Baseline 1’s transformer structure results in slower processing as the input sequence grows.
@@
 Based on the results, we draw the following two conclusions.
-First, our C-VUE demonstrates the best continuous perception ability (Capability 4).
+First, our CARS demonstrates the best continuous perception ability (Capability 2).
 It surpasses batch processing methods ranging from 0.5 to 10.7.
 Its accuracy is twice that of traditional LSTM-based baseline methods.
-Second, our approach can effectively adapt to events of varying scales (Capability 1).
-In the hours-long video, events lasted from tens of frames to hundreds of frames.
-Our C-VUE achieves the best accuracy, which outperforms the current sliding-window and state-assist methods.
-5.4. Further Evaluation of Capability 5
-We further profile the memorization abilities on three multi-label Finegym datasets to evaluate our memorization capability.
-The k value represents how many events are concatenated to form a testing data sample.
-Therefore, a higher k value indicates a longer video with more labels.
-All baseline methods use the model trained on the FineGym99 dataset.
-Baselines 1 and 2 are evaluated with input strategy 1, where they can achieve the best accuracy.
-We utilize three metrics.
-The Jaccard Index [^1] The threshold ‘th’ signifies that in the model’s output vector, values greater than ‘th’ are considered as detecting a particular category, otherwise not recognized.
-is calculated as the intersection of predicted results and ground truth divided by their union.
-We report the results in Table 3 and draw the following conclusions.
-Our C-VUE demonstrates an superior memorization capabilities (Capability 5).
-Across all configurations and metrics, our method surpasses the baseline approaches.
-Our C-VUE excels at tasks involving frame-by-frame understanding.
-Surprisingly, baseline methods that process frames in batches (Baselines 1 and 2), incorporating all frames at each time step, exhibit notably lower accuracy compared to streaming methods (Ours and Baseline 3).
-This difference arises from the fact that streaming methods prioritize the present frame while retaining information from past frames in streaming data processing.
-Table 3. Evaluation results on three Multi-label FineGym@k datasets.
-Metrics Methods k=1 k=3 k=5
-Jaccard Index (th=0.5) Baseline 1 0.0351 0.0172 0.0087
-Baseline 2 0.0194 0.0281 0.0313
-Baseline 3 0.0906 0.0484 0.0287
-Ours 0.1210 0.0578 0.0341
-Jaccard Index (th=0.3) Baseline 1 0.0803 0.04396 0.0228
-Baseline 2 0.0101 0.0152 0.0200
-Baseline 3 0.1426 0.0880 0.0554
-Ours 0.1752 0.0957 0.0604
-Jaccard Index (th=0.1) Baseline 1 0.1324 0.0802 0.0441
-Baseline 2 0.0101 0.0152 0.0200
-Baseline 3 0.1428 0.1254 0.0935
-Ours 0.1846 0.1334 0.0936
-mAP Baseline 1 0.1590 0.1196 0.0905
-Baseline 2 0.0465 0.0374 0.0353
-Baseline 3 0.2338 0.1645 0.1277
-Ours 0.2678 0.1739 0.1379
-ranking AP Baseline 1 0.3879 0.3296 0.2771
-Baseline 2 0.4533 0.4338 0.4108
-Baseline 3 0.5856 0.4906 0.4250
-Ours 0.5894 0.4916 0.4346
-Figure 9. Processing speed of understanding a video with 10k frames on different platforms.
-5.5. Real-time Inference on Edge
+Second, our approach can effectively adapt to activities of varying scales (Capability 1).
+In the hours-long video, activities lasted from tens of frames to hundreds of frames.
+Our CARS achieves the best accuracy, which outperforms the current sliding-window and state-assist methods.
+4.4 Real-time Inference on Edge
 We measure the system performance on three mobile platforms with two long video clips.
-The long video clips were excerpted from the All-event-in-one FineGym dataset, one containing 10k frames and the other 100k frames.
+The long video clips were excerpted from the All-activity-in-one FineGym dataset, one containing 10k frames and the other 100k frames.
 We conducted each experiment three times and calculated the average.
@@
 Our experimental findings are as follows.
-Understanding the video stream in real time.
-In Figure 9, it can be observed that our method achieves real-time processing (¿30fps) on different platforms.
+Recognizing the video stream in real time.
+In Figure 9, it can be observed that our method achieves real-time processing ($>$30fps) on different platforms.
 Even on Macbook Air and Jetson, our speed is 2 to 20 times faster than the baseline methods.
@@
 When processing a 100k frame video, Baseline 1 drops to a speed of 9fps.
-6. Case Study: Video Encoder for LLM-based Video Foundation Models
-Figure 10. Applying our history modeling into VideoLLaMa2 as a video encoder.
-In this section, we discuss how the design of our system, C-VUE, improves video foundation models.
-As a case study, we analyze one of the SOTA video foundation models, VideoLLaMa2 (damonlpsg2024videollama2, ) (left side of Figure 10).
+5 Case Study: Video Encoder for Large Video Models
+Figure 10: Applying our CARS into VideoLLaMa2 as a video encoder.
+In this section, we discuss how the design of our system, CARS, improves large video models in video activity recognition and understanding.
+As a case study, we analyze one of the SOTA large video models, VideoLLaMa2 [8] (left side of Figure 10).
 During video processing, VideoLLaMa2 first samples $N$ frames from the video.
@@
 Simultaneously, the user’s text prompt is tokenized, combined with the video tokens, and input into the LLM for inference.
-6.1. Problem Analysis
+5.1 Problem Analysis
 VideoLLaMa2 introduces an STC projector as the projection layer.
@@
 Regardless of the video length or event scale, this fixed frame sampling inevitably reduces the model’s accuracy.
-6.2. Our design
-We adopt the design of our C-VUE to build a new projection layer, named the C-VUE projector, which replaces the STC projector.
-This new projection utilizes the long-range history modeling to enable a high sampling rate in a remarkably efficient manner.
+5.2 Our design
+We adopt the design of our CARS to build a new projection layer, named the CARS projector, which replaces the STC projector.
+This new projection utilizes our activity-aware state update to maintain the video context efficiently.
+Our state-assisted design can efficiently increase the sampling rate of the input video, far exceeding the existing designs with 8 or 16 frames, without increasing the processing time.
 Figure 10 (right) illustrates our design.
 When a video frame is input, the projector first processes the spatial information within the frame by applying mean pooling to patch embeddings.
-Next, the history modeling layer processes the history information, capturing relationships over time.
-Its processing speed remains consistent regardless of video length, making it more efficient.
-Table 4. Evaluation results on the VideoLLaMa2.
+Next, our state-assistant module processes the input video, capturing relationships over time.
+Table 3: Evaluation results on the VideoLLaMa2.
 VideoChatGPT
@@
 Ours 2.41 39.65 48.6
-6.3. Experiments
+5.3 Experiments
 Implementation.
-Using the official videollama2 codebase (video_llama_impl, ), we implemented our C-VUE projector.
-This projector consists of mean pooling, a linear layer with ReLU, and the history modeling layer of our C-VUE with a code dimension of 2048, followed by another linear layer with ReLU.
+Using the official videollama2 codebase [6], we implemented our CARS projector.
+This projector consists of mean pooling, a linear layer with ReLU, and the history modeling layer of our CARS with a code dimension of 2048, followed by another linear layer with ReLU.
 The total parameter count of our projector is 450 M.
@@
 During fine-tuning, only the large language model (LLM) and the projection layer were trained.
-We used the videoChatGPT dataset (maaz2023video, ) for fine-tuning both the original model structure and our proposed model structure.
+We used the videoChatGPT dataset [25] for fine-tuning both the original model structure and our proposed model structure.
 All experiments were performed on a machine with eight A600 GPUs.
@@
 Results.
-Testing was conducted on the videoChatGPT dataset (in-distribution), ActivityNet-QA (yu2019activityqa, ) (zero-shot), and MSVD-QA (xu2017video, ) (zero-shot), and the results are reported in Table 4.
+Testing was conducted on the videoChatGPT dataset (in-distribution), ActivityNet-QA [44] (zero-shot), and MSVD-QA [43] (zero-shot), and the results are reported in Table 3.
 For the VideoChatGPT dataset, correctness was measured on a 5-point scale.
@@
 On the other two zero-shot datasets, our approach achieved improvements of 1.19% and 4%, respectively.
-These results demonstrate that our projector design effectively improves model accuracy, highlighting a new design concept for video projection for video foundation models.
-7. Related Work
-Video Analytics & Understanding.
-Video analytics and understanding involve the automatic extraction of valuable information from video including classification (wu2020emo, ), detection (tong2021large, ; jiang2021flexible, ), tracking (voigtlaender2019mots, ), identification (yi2020eagleeye, ), and privacy protection (wu2021pecam, ).
-Many approaches treat continuous video analysis as a task of image understanding on video frames.
-Adaptive frame sampling (hwang2022cova, ; moll2022exsample, ), cooperation between small and large models (liu2020continuous, ), and collaboration across multiple computation resources (chen2015glimpse, ; liu2019edge, ; damonlpsg2024videollama2, ) are proposed to enable efficient video analytics on video frames.
-To analyze the events that involve multiple frames, 3D convolution is used to perform the video understanding (slowfast2019fei, ).
-However, this approach can cause significant system performance issues (zhu2020comprehensive, ).
-To improve efficiency, MoviNet (kondratyuk2021movinets, ) creates efficient 3D convolution architectures.
-Many efforts also explore temporal modeling without 3D convolution (Lin_2019_ICCV, ; Jiang_2019_ICCV, ; Li_2020_CVPR, ).
-MViTv2 (li2022mvitv2, ) uses improved pooling attention for spatiotemporal recognition to improve the attention’s efficiency.
-These methods improve video understanding by focusing on relationships in a short video clip.
-They still struggle with system performance issues in long-form and streaming videos.
-Video Foundation Models.
-Recent studies have explored building video foundation models using large language models (LLMs).
-These models achieve impressive results in multiple video understanding tasks (madan2024foundation, ).
-Methods like VideoLaVIT (jin2024video, ), VideoLLaVA (lin2023video, ), Video-LLaMA (zhang2023video, ), and Video-LLaMA2 (damonlpsg2024videollama2, ) use multi-modal pre-training to align visual information with LLMs through projection layers.
-VideoChatGPT (maaz2023video, ) improves video-based conversations by integrating visual encoders with LLMs and instruction tuning.
-The main challenge in video understanding is handling memory requirements and model capacity, especially for long videos.
-LaViLa (zhao2023learning, ) addresses this by turning LLMs into narrators, allowing them to generate detailed descriptions of long videos from visual inputs.
-ChatVideo (wang2023chatvideo, ) and MM-VID (lin2023mm, ) convert videos into text for better comprehension.
-However, these methods still struggle to model long-range temporal information in streaming videos.
-Alternatives to Foundation Models for Video Understanding.
-The series of works on Mamba have gained significant attention for overcoming the high computational complexity of Transformers as LLMs.
-Recently, Mamba has been applied to various computer vision tasks, showing early progress (liu2024vision, ).
-However, its application is still in its infancy.
-Mamba-ND (li2024mambandselectivestatespace, ) is the first to demonstrate Mamba can handle high-dimensional data and experimentally show the effectiveness in video action recognition.
-VideoMamba (li2024videomambastatespacemodel, ) builds an SSM-based model for video understanding.
-It shares a similar design with VideoLlama (damonlpsg2024videollama2, ) but replaces the Transformer-based LLM with a Mamba-based one for efficiency purposes.
-VideoMamba Suite (chen2024video, ) further evaluates Mamba’s performance by integrating it into existing video understanding architectures, showing its potential to handle complex spatial-temporal dynamics.
-Unlike these works that use “all frames at once” or “fixed window size” strategies for Mamba and Transformer video understanding, we propose a lightweight continuous video understanding design with each frame as the input to the state-assistance module.
-8. Discussion & Conclusion
-Discussion.
-Our work presents a solution for continuous video understanding, but there are still unexplored issues we hope to address with the community in the future.
-The memorization upper bound of the state modeling approach has not been fully explored.
-Although several improvements were made, a fixed-size hidden state will fail to model infinitely long videos.
-Future work should figure out the modeling capacity of such methods.
-Additionally, the generalization ability of these approaches to more and more tasks remains unclear, which is crucial for designing new video foundation models based on this design.
-Conclusion.
-Continuous video understanding is important for complex multimodal human-computer interaction.
+These results demonstrate that our projector design effectively improves model accuracy, highlighting a new design concept for video projection for large video models.
+6 Conclusion
+Continuous video activity recognition in streaming video is important for complex multimodal human-computer interaction, e.g., embodied AI.
 However, achieving this on resource-limited edge devices is challenging.
-This paper proposes C-VUE, an adaptive state modeling approach, for continuous video understanding.
-Experimental results show that C-VUE can achieve real-time continuous video understanding on typical edge platforms.
-As a case study, we apply C-VUE’s design to improve video foundation models and obtain a large accuracy improvement.
+This paper proposes CARS, a state-assistant approach for continuous video activity recognition in streaming video.
+It consists of two key designs: activity spatial feature extraction to reduce spatial redundancy and activity-aware state update to maintain the video context effectively.
+Experimental results show that CARS can achieve real-time continuous video recognition on typical edge platforms.
+We believe that our CARS offers a novel idea for future streaming video processing.